### v1

Abstract
I consider treatment-effect estimation with a panel data with two periods, using a first difference regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. A design-based justification of this estimator assumes that $\Delta D_{g}$ is as good as randomly assigned. This note shows that if one posits a causal model in levels between the treatment and the outcome, then $\Delta D_{g}$ randomly assigned implies that $\Delta D_{g}$ is independent of $D_{g,1}$. This is a very strong, fully testable condition. If $\Delta D_{g}$ is not independent of $D_{g,1}$, the first-difference estimator can still identify a convex combination of effects if $\Delta D_{g}$ is randomly assigned and treatment effects do not change over time, or if the the treatment path $(D_{g,1},D_{g,2})$ is randomly assigned and either of the following conditions hold: i) $D_{g,1}$ and $D_{g,2}$ are binary; ii) $D_{g,1}$ and $D_{g,2}$ have the same variance; iii) $D_{g,1}$ and $D_{g,2}$ are uncorrelated or negatively correlated. I use these results to revisit acemoglu2016import.
1 Introduction
To estimate a treatment’s effect with a two-periods panel data set, researchers often use a first-difference regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. Let $\widehat{\beta}$ denote the coefficient on $\Delta D_{g}$. One may invoke a model-based parallel trends assumption to justify $\widehat{\beta}$. However, under parallel trends, $\widehat{\beta}$ may not estimate a convex combination of treatment effects, outside of the canonical design where all units are untreated at period one and treatment is binary (dcDH2020; callaway2021difference). An alternative, design-based justification of $\widehat{\beta}$ assumes that $\Delta D_{g}$ is as good as randomly assigned. This note considers this second justification.
If one posits a causal model in first difference, where different values of the treatment evolution $\Delta D_{g}$ may give rise to different counterfactual values of the outcome evolution $\Delta Y_{g}(\Delta d)$, then it is well-known that $\widehat{\beta}$ estimates a well-defined causal effect if $\Delta D_{g}$ is randomly assigned, even if treatment effects are heterogeneous across units. However, the first-difference specification does not always come from the researcher’s belief that the causal relation between the treatment and the outcome only arises in first difference. Rather, the first-difference specification may sometimes come from the researcher’s belief that first-differencing is necessary to net out time-invariant counfounders.
If one instead posits a causal model in levels, where different values of the treatment $D_{g,t}$ may give rise to different counterfactual values of the outcome $Y_{g,t}(d)$, I show that $\Delta D_{g}$ can be as good as randomly assigned “essentially only if” [^1] I borrow the “essentially only if” terminology from roth2023parallel. $\Delta D_{g}$ is independent of $D_{g,1}$. This is a very strong condition. For instance, it cannot hold if $D_{g,1}$ and $D_{g,2}$ have the same bounded support. This condition is also fully testable, as both $\Delta D_{g}$ and $D_{g,1}$ are observed. A simple way to test it is for instance to regress $\Delta D_{g}$ on $D_{g,1}$.
Then, I show that if one is ready to assume that the treatment effect does not change over time, one can have a causal model in levels and a randomly-assigned $\Delta D_{g}$ even if $\Delta D_{g}$ is not independent of $D_{g,1}$.
Finally, I show that under a causal model in levels with time-varying effects, if $\Delta D_{g}$ is not independent of $D_{g,1}$ and if the treatments $D_{g,1}$ and $D_{g,2}$ are non-binary, positively correlated, and have different variances, then $\widehat{\beta}$ may identify a non-convex combination of time-varying effects, even if the treatment paths $(D_{g,1},D_{g,2})$ are randomly assigned, which is stronger than assuming that $\Delta D_{g}$ is randomly assigned. Intuitively, under a causal model in levels with time-varying effects, the first-difference regression is misspecified.
As an empirical application, I revisit acemoglu2016import, who run a first-difference regression with three periods and two first-differences, to estimate the effect of US industries’ Chinese-imports penetration ratio on their employment. To fit in the simple setting considered in this note, I reestimate their regression, keeping only two time periods, 1999 and 2007, and one first difference. Doing so, I obtain a coefficient very close to that in the paper. I start by regressing $\Delta D_{g}$ on $D_{g,1}$, and I find that the two variables are strongly positively correlated: industries where the import-penetration ratio increased the most already had a larger import-penetration ratio in 1999. Then, I show that under a causal model in levels with time-varying effects, if $\Delta D_{g}$ is not independent of $D_{g,1}$, $\widehat{\beta}$ estimates a non-convex combination of the effects of the Chinese-import-penetration ratio in 1999 and 2007, where the average 1999 effect receives a weight equal to $-0.30$. This is because the Chinese-import-penetration ratios are non-binary, positively correlated treatment variables, with a much larger variance in 2007 than in 1999.
Some of the ideas in this note originated while working on de2021more, see in particular Sections 2.1 and 3.3 therein. As noted page 17 therein, these ideas apply more generally than to the first-difference 2SLS Bartik regressions considered by de2021more. Accordingly, these ideas are now presented in this standalone note, and future versions of de2021more will not contain material overlapping with that in this note.
2 Theoretical Results
2.1 Set-up
Two-periods panel.
One is interested in estimating the effect of a treatment on an outcome, using a panel data set, with $G$ independent and identically distributed units indexed by $g$, and with two time periods. For $t\in\{1,2\}$, let $Y_{g,t}$ and $D_{g,t}$ respectively denote the outcome and treatment of $g$ at $t$.
Assumption 1
$(Y_{g,1},Y_{g,2},D_{g,1},D_{g,2})_{g=1,...,G}$ are i.i.d.
As units are assumed i.i.d., I sometimes drop the $g$ subscript below.
First-difference regression.
Let $\Delta$ denote the first difference operator. To estimate the treatment’s effect, researchers often regress $\Delta Y_{g}$ on an intercept and $\Delta D_{g}$, leading to
$\widehat{\beta}:=\frac{\sum_{g}(\Delta D_{g}-\Delta D_{.})\Delta Y_{g}}{\sum_{g}(\Delta D_{g}-\Delta D_{.})^{2}},$
where $\Delta D_{.}=1/G\sum_{g}\Delta D_{g}.$ Assuming that $(\Delta Y,\Delta D)$ has a finite second moment, under Assumption 1 the probability limit of $\widehat{\beta}$ when $G\rightarrow+\infty$ is equal to
$\beta:=\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}.$
2.2 Under a causal model in first-difference, $\beta$ identifies an average slope if the first-differenced treatment is as good as randomly assigned.
A linear causal model in first difference.
In this section, I consider a causal model in first-difference, where different values of the treatment evolution can lead to different counterfactual outcome evolutions. Accordingly, let $\Delta Y_{g}(\Delta d)$ denote $g$’s potential first-differenced outcome if $\Delta D_{g}=\Delta d$. $\Delta Y_{g}(0)$ is $g$’s potential first-differenced outcome if $\Delta D_{g}=0$, i.e. if $D_{g,2}=D_{g,1}$, namely in the counterfactual where $g$’s treatment does not change. I assume that the treatment evolution has a linear effect on the outcome evolution.
Assumption 2
$\Delta Y(\Delta d)=\Delta Y(0)+S_{\delta}\times\Delta d.$ (1)
Assumption 2 is a linear effect assumption, but it allows for heterogeneous effects across units: the slope of the potential first-differenced outcome function $S_{\delta}$ is a random variable, that may vary across units. Under Assumption 2,
$\Delta Y=\Delta Y(0)+S_{\delta}\times\Delta D.$ (2)
As good as randomly assigned first-differenced treatment.
Assumption 3
$\Delta D\perp\!\!\!\perp(\Delta Y(0),S_{\delta}).$ (3)
Assumption 3 requires that the first-differenced treatment be independent of units’ outcome evolutions without a treatment change and of their treatment effect.
Theorem 1
If Assumptions 2 and 3 hold, $\beta=E(S_{\delta})$.
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,\Delta Y(0))+cov(\Delta D,S_{\delta}\times\Delta D)}{V(\Delta D)}$
$=$ $\frac{E(S_{\delta}\times\Delta D^{2})-E(\Delta D)E(S_{\delta}\times\Delta D)}{V(\Delta D)}$
$=$ $E(S_{\delta}).$
The first equality follows from (2). The second and third equalities follow from Assumption 3. QED.
Theorem 1 is a well-known result. It shows that $\beta$ identifies the average slope $E(S_{\delta})$ under Assumptions 2 and 3. Under Assumption 3, $\beta$ can still identify a well-defined causal effect if the treatment effect is non-linear: the linearity assumption is made to simplify the discussion.
2.3 Unpacking the implications of Assumptions 2 and 3 under a causal model in levels.
A linear causal model in levels.
If one is ready to posit that different counterfactual treatment evolutions can lead to different outcome evolutions, one may also be ready to posit that different counterfactual treatment levels can lead to different outcome levels. Then, let $Y_{g,t}(d)$ denote $g$’s potential outcome if $D_{g,t}=d$. $Y_{g,t}(0)$ is $g$’s potential outcome at $t$ if $D_{g,t}=0$, namely without treatment. Implicitly, our potential outcome notation rules out anticipation effects where $D_{g,2}$ could affect the period-one outcome, and dynamic effects where $D_{g,1}$ could affect the period-two outcome. To simplify, let us again assume that the model in levels is linear.
Assumption 4
For $t\in\{1,2\}$,
$Y_{t}(d)=Y_{t}(0)+S_{t}d.$ (4)
The slope $S_{t}$ is indexed by $t$, as the treatment’s effect might be time-varying.
Restating Assumption 3 under the causal models in levels and in first difference.
Under the linear causal model in levels in Assumption 4,
$\Delta Y(0)=Y_{2}(D_{1})-Y_{1}(D_{1})=Y_{2}(0)-Y_{1}(0)+(S_{2}-S_{1})D_{1}.$
Then, combining the linear causal models in first-difference and in levels yields
$S_{\delta}=S_{2}.$
Thus, Assumption 3 is equivalent to
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0)+(S_{2}-S_{1})D_{1},S_{2}).$ (5)
A test of a sufficient condition for Assumption 3, under the causal models in levels and in first difference.
In view of (5), a sufficient, more interpretable sufficient condition for Assumption 3 to hold is
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2},D_{1}).$ (6)
As $\Delta D$ and $D_{1}$ are observed, (6) has the following, fully testable implication:
$\Delta D\perp\!\!\!\perp D_{1}.$ (7)
(7) is a strong condition. For instance, it cannot hold if $D_{1}$ and $D_{2}$ have the same bounded support. A simple way to test (7) is to run a linear regression of $\Delta D$ on $D_{1}$, though this test will not be consistent against all alternatives. As (7) is a testable implication of a sufficient condition for Assumption 3, Assumption 3 may still hold when (7) is rejected. Still, it may be difficult to convincingly argue for Assumption 3 when (7) is rejected.
If one is ready to assume constant effects over time, Assumption 3 no longer implies strong restrictions.
In view of (5), if one is ready to assume that $S_{2}=S_{1}$, Assumption 3 is equivalent to
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{2}),$
an untestable condition very close in spirit to the idea that $\Delta D$ is as good as randomly assigned. Therefore, it is only if treatment effects are time-varying that Assumption 3 implies strong restrictions when combined with a causal model in levels.
2.4 Under a causal model in levels, if $\Delta D$ is not independent of $D_{1}$ and treatment effects change over time, $\beta$ may not identify a convex combination of effects.
In this section, I consider the following assumption:
Assumption 5
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{2}),$ (8)
and
$(D_{1},D_{2})\perp\!\!\!\perp\Delta S.$ (9)
(8) is almost equivalent to (6), except that the testable and potentially problematic part of (6), involving the independence between $\Delta D$ and $D_{1}$, has been taken out. (9) requires that units’ treatment path $(D_{1},D_{2})$ be independent of the change in their treatment effect $\Delta S$. Thus, Assumption 5 is somewhere in between assuming that the treatment change $\Delta D$ is as good as randomly assigned, and assuming that the treatment path $(D_{1},D_{2})$ is as good as randomly assigned, which is stronger.
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})$ (10)
or the even stronger condition
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{1}(0),Y_{2}(0),S_{1},S_{2})$ (11)
are both sufficient for Assumption 5 to hold. Under (11), estimating a first-difference regression is not necessary anymore: to estimate the treatment’s effect, one can merely regress $Y_{1}$ on an intercept and $D_{1}$ to estimate $E(S_{1})$, and one can regress $Y_{2}$ on an intercept and $D_{2}$ to estimate $E(S_{2})$.
Theorem 2
If Assumptions 4 and 5 hold,
$\beta=\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\Delta D)+cov(\Delta D,\Delta SD_{1})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}$
$=$ $\frac{V(\Delta D)E(S_{2})+cov(\Delta D,D_{1})E(\Delta S)}{\sum_{t^{\prime}=1}^{2}V(D_{t^{\prime}})-cov(D_{1},D_{2})}$
$=$ $\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
The first equality follows from Assumption 4. The second follows from Assumption 5. QED.
Theorem 2 shows that under a linear causal model in levels and Assumption 5, $\beta$ estimates a weighted sum of the average slopes $E(S_{1})$ and $E(S_{2})$. Weights are guaranteed to be positive if the treatment $D_{g,t}$ is binary. [^2] To see this, note that for binary variables, $E(D_{1}D_{2})\leq\min(E(D_{1}),E(D_{2}))$. Therefore, $V(D_{t})-cov(D_{1},D_{2})\geq V(D_{t})-\min(E(D_{1}),E(D_{2}))+E(D_{1})E(D_{2})\geq 0$. This is consistent with the results of athey2022design and arkhangelsky2021double, who have shown that with a binary randomized treatment, OLS TWFE regressions always estimate a convex combination of effects. By Cauchy-Schwarz’s inequality, weights are also guaranteed to be positive if $V(D_{1})=V(D_{2})$. Finally, weights are also guaranteed to be positive if $cov(D_{1},D_{2})=0$, meaning that the treatments are not serially correlated, or if $cov(D_{1},D_{2})\leq 0$, meaning that the treatments are negatively correlated. On the other hand, weights could be negative if $D_{1}$ and $D_{2}$ are non-binary, positively correlated, and have different variances.
Importantly, replacing Assumption 5 by the stronger condition in (10) or by the even stronger condition in (11) does not lead to a different decomposition from that in Theorem 2: even with a randomly assigned treatment path, a first-difference regression may not estimate a convex combination of effects with positively correlated non-binary treatments, under a causal model in levels with time-varying effects. This is simply due to the fact that with a causal model in levels with time-varying effects, the first-difference regression is misspecified. Under (10),
$E(\Delta Y|D_{1},D_{2})=E(Y_{2}(0)-Y_{1}(0))+\Delta DE(S_{2})+D_{1}(E(S_{2})-E(S_{1})).$
To correct for this misspecification, one can just regress $\Delta Y$ on an intercept, $\Delta D$, and $D_{1}$.
3 Application
Data.
I use the industry-level panel data set constructed by acemoglu2016import. In their Table 2 Column (2), the authors estimate a first-difference regression of industries’ employment evolutions on changes in industries’ Chinese-import-penetration ratio. The regression leverages three time periods and two first differences. To fit in the two-period setting considered in this paper, I estimate a similar regression, with the same weights, clustering standard errors at the same level as the authors, but keeping only two time periods and one first difference, from 1999 to 2007. Doing so, I find that $\widehat{\beta}=-0.78$ (s.e.=$0.22$), which is very close to the coefficient in the paper’s Table 2 Column (2) ($-0.81$, s.e.=$0.16$). According to this regression, an increase in the Chinese-import-penetration ratio decreases US employment.
Test of (7).
I start by testing (7), by regressing $\Delta D$, the change in industries’ import-penetration ratio, on $D_{1}$, industries’ import-penetration ratio in 1999. The coefficient on $D_{1}$ is equal to $0.74$, and it is highly significant (s.e.=$0.16$). Thus, the industries that experienced the largest growth of their imports penetration ratio from 1999 to 2007 already had a larger imports penetration ratio in 1999.
Estimating the weights attached to $\beta$ in Theorem 2.
Next, I estimate the weights attached to $\beta$ in Theorem 2, namely under a causal model in levels with time varying effects, and without assuming that $\Delta D$ is independent of $D_{1}$. Doing so, I find that $\beta$ estimates a weighted sum of the average effects of import penetration in 1999 and 2007, where the 2007 average effect receives a weight equal to 1.3, while the 1999 average effect receives a weight equal to -0.3. Thus, under a causal model in levels and Assumption 5, the first-difference regression does not estimate a convex combination of effects, if the effect of import penetration on US employment changes from 1999 to 2007.
Regressing $\Delta Y$ on $\Delta D$ and $D_{1}$.
Under (10), one can estimate the 2007 average effect of import penetration, as well as the difference between the 2007 and 1999 average effect, by regressing $\Delta Y$ on $\Delta D$ and $D_{1}$. The coefficient on $\Delta D$ is equal to $-0.64$ (s.e.=0.24), which is 18% smaller than $\widehat{\beta}$. The coefficient on $D_{1}$ is equal to $-0.45$ and it is marginally significant (s.e. =$0.25$), thus suggesting that the effect of import penetration has become more negative between 1999 and 2007. Note however that the regression of $\Delta Y$ on $\Delta D$ and $D_{1}$ relies on (10). This condition requires that industries’ import penetration paths $(D_{1},D_{2})$ be randomly assigned, which is stronger than assuming that changes in import penetration ratios $\Delta D$ are randomly assigned.