### v2

Abstract
I consider treatment-effect estimation with a two-periods panel, using a first-difference regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. To justify this regression, one may assume that $\Delta D_{g}$ is as good as randomly assigned, namely uncorrelated to the residual of the first-differenced model and to the treatment’s effect. This note shows that if one posits a causal model in levels between the treatment and the outcome, then the residual of the first-differenced model is a function of $D_{g,1}$, so $\Delta D_{g}$ uncorrelated to that residual essentially implies that $\Delta D_{g}$ is uncorrelated to $D_{g,1}$. This is a strong, testable condition. If $\Delta D_{g}$ is correlated to $D_{g,1}$, assuming that $\Delta D_{g}$ is uncorrelated to the treatment effect and to the remaining terms of the residual may not be sufficient to have that the first-difference regression identifies a convex combination of treatment effects. I use these results to revisit Acemoglu et al. (2016).
1 Introduction
To estimate a treatment’s effect with a two-periods panel data set, researchers often use a first-difference regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. Let $\widehat{\beta}$ denote the coefficient on $\Delta D_{g}$. One may invoke a model-based parallel trends assumption to justify $\widehat{\beta}$. However, under parallel trends, $\widehat{\beta}$ may not estimate a convex combination of treatment effects, outside of the canonical design where all units are untreated at period one and treatment is binary (de Chaisemartin and D’Haultfœuille, 2020; Callaway et al., 2021). An alternative, design-based justification of $\widehat{\beta}$ assumes that $\Delta D_{g}$ is as good as randomly assigned. This note considers this second justification.
Assume that one posits a causal model in first difference, where different values of the treatment evolution $\Delta D_{g}$ may give rise to different counterfactual values of the outcome evolution $\Delta Y_{g}(\Delta d)$. In this model, it is well-known that if $\Delta D_{g}$ is as-good-as randomly assigned, namely uncorrelated to the residual of the first-differenced model and to the effect of the first-differenced treatment, then $\widehat{\beta}$ estimates an average effect of the first-differenced treatment on the first-differenced outcome. However, the first-difference specification does not always come from the researcher’s belief that the causal relation between the treatment and the outcome only arises in first difference. Rather, the first-difference specification may sometimes come from the researcher’s belief that first-differencing is necessary to net out time-invariant counfounders.
If one instead posits a causal model in levels, where different values of the treatment $D_{g,t}$ may give rise to different counterfactual values of the outcome $Y_{g,t}(d)$, I show that the residual of the first-differenced model is a function of $D_{g,1}$, so $\Delta D_{g}$ uncorrelated to that residual essentially implies that $\Delta D_{g}$ is uncorrelated to $D_{g,1}$. This is a strong condition, which can be tested by regressing $\Delta D_{g}$ on $D_{g,1}$.
Then, I show that if one is ready to assume that the treatment effect does not change over time, the residual of the first-differenced model is no longer a function of $D_{g,1}$, so one can have that $\Delta D_{g}$ is uncorrelated to that residual even if $\Delta D_{g}$ is correlated to $D_{g,1}$.
Finally, if $\Delta D_{g}$ is correlated to $D_{g,1}$ but one is not ready to assume that the treatment effect does not change over time, assuming that $\Delta D_{g}$ is uncorrelated to the treatment effect and to the remaining terms of the residual may not be sufficient to have that the first-difference regression identifies a convex combination of treatment effects. Specifically, if the treatments $D_{g,1}$ and $D_{g,2}$ are non-binary, positively correlated, and have different variances, then $\widehat{\beta}$ may identify a non-convex combination of time-varying effects, even if the treatment paths $(D_{g,1},D_{g,2})$ are randomly assigned, which is stronger than assuming that $\Delta D_{g}$ is randomly assigned.
As an empirical application, I revisit Acemoglu et al. (2016). One of their regressions is a first-difference regression with three periods and two first-differences, to estimate the effect of US industries’ Chinese-imports-penetration ratio on their employment. To fit in the simple setting considered in this note, I reestimate their regression, keeping only two time periods, 1999 and 2007, and one first difference. Doing so, I obtain a coefficient very close to that in the paper. I start by regressing $\Delta D_{g}$ on $D_{g,1}$, and I find that the two variables are strongly positively correlated: industries where the import-penetration ratio increased the most already had a larger import-penetration ratio in 1999. Then, I show that under a causal model in levels with time-varying effects, $\widehat{\beta}$ estimates a non-convex combination of the effects of the Chinese-import-penetration ratio in 1999 and 2007, where the average 1999 effect receives a weight equal to $-0.30$. This is because the Chinese-import-penetration ratios are non-binary, positively correlated treatment variables, with a much larger variance in 2007 than in 1999.
Some of the ideas in this note originated while working on de Chaisemartin and Lei (2021), see in particular Sections 2.1 and 3.3 therein. That paper considers first-difference 2SLS Bartik regressions, that are sometimes justified by assuming that the first-difference shocks used to construct the Bartik instrument are as good as randomly assigned (Borusyak et al., 2022). As noted in de Chaisemartin and Lei (2021), issues with design-based justifications of first-difference regressions are not specific to 2SLS Bartik regressions, which is why these ideas are now presented in this standalone note. Future versions of de Chaisemartin and Lei (2021) will not contain material overlapping with that in this note.
2 Theoretical Results
2.1 Set-up
Two-periods panel.
One is interested in estimating the effect of a treatment on an outcome, using a panel data set, with $G$ independent and identically distributed units indexed by $g$, and with two time periods. For $t\in\{1,2\}$, let $Y_{g,t}$ and $D_{g,t}$ respectively denote the outcome and treatment of $g$ at $t$.
Assumption 1
$(Y_{g,1},Y_{g,2},D_{g,1},D_{g,2})_{g=1,...,G}$ are i.i.d.
As units are assumed i.i.d., I sometimes drop the $g$ subscript below.
First-difference regression.
Let $\Delta$ denote the first difference operator. To estimate the treatment’s effect, researchers often regress $\Delta Y_{g}$ on an intercept and $\Delta D_{g}$, leading to
$\widehat{\beta}:=\frac{\sum_{g}(\Delta D_{g}-\Delta D_{.})\Delta Y_{g}}{\sum_{g}(\Delta D_{g}-\Delta D_{.})^{2}},$
where $\Delta D_{.}=1/G\sum_{g}\Delta D_{g}.$ Assuming that $(\Delta Y,\Delta D)$ has a finite second moment, under Assumption 1 the probability limit of $\widehat{\beta}$ when $G\rightarrow+\infty$ is equal to
$\beta:=\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}.$
2.2 Under a linear causal model in first-difference, $\beta$ identifies an average slope if the first-differenced treatment is as good as randomly assigned.
A linear causal model in first difference.
In this section, I consider a causal model in first-difference, where different values of the treatment evolution can lead to different counterfactual outcome evolutions. Accordingly, let $\Delta Y_{g}(\Delta d)$ denote $g$’s potential first-differenced outcome if $\Delta D_{g}=\Delta d$. $\Delta Y_{g}(0)$ is $g$’s potential first-differenced outcome if $\Delta D_{g}=0\Leftrightarrow D_{g,2}=D_{g,1}$, meaning that $g$’s treatment does not change. I assume that the treatment evolution has a linear effect on the outcome evolution.
Assumption 2
$\Delta Y(\Delta d)=\Delta Y(0)+S_{\delta}\times\Delta d.$ (1)
Assumption 2 is a linear effect assumption, but it allows for heterogeneous effects across units: the slope of the potential first-differenced outcome function $S_{\delta}$ is a random variable, that may vary across units. Under Assumption 2,
$\Delta Y=\Delta Y(0)+S_{\delta}\times\Delta D.$ (2)
As good as randomly assigned first-differenced treatment.
Assumption 3
$cov(\Delta D,\Delta Y(0))=cov(\Delta D,S_{\delta})=cov\left(\left(\Delta D\right)^{2},S_{\delta}\right)=0.$ (3)
Assumption 3 requires that the first-differenced treatment be uncorrelated to units’ outcome evolutions without a treatment change, the residual in the model relating the first-differenced outcome and the first-differenced treatment. Assumption 3 also requires that the first-differenced treatment and its square be uncorrelated to the slope $S_{\delta}$.
Theorem 1
If Assumptions 2 and 3 hold, $\beta=E(S_{\delta})$.
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,\Delta Y(0))+cov(\Delta D,S_{\delta}\times\Delta D)}{V(\Delta D)}$
$=$ $\frac{E\left(S_{\delta}\times\left(\Delta D\right)^{2}\right)-E(\Delta D)E(S_{\delta}\times\Delta D)}{V(\Delta D)}$
$=$ $E(S_{\delta}).$
The first equality follows from (2). The second and third equalities follow from Assumption 3. QED.
Theorem 1 is a well-known result. It shows that $\beta$ identifies the average slope $E(S_{\delta})$ under Assumptions 2 and 3. Under Assumption 3, $\beta$ can still identify a well-defined causal effect if the treatment effect is non-linear: the linearity assumption is made to simplify the discussion.
2.3 Unpacking Assumption 3 under a causal model in levels.
A linear causal model in levels.
Assume that instead of positing that different counterfactual treatment evolutions can lead to different outcome evolutions, one prefers to posit that different counterfactual treatment levels can lead to different outcome levels. Then, let $Y_{g,t}(d)$ denote $g$’s potential outcome if $D_{g,t}=d$. $Y_{g,t}(0)$ is $g$’s potential outcome at $t$ if $D_{g,t}=0$, namely without treatment. [^1] Implicitly, our potential outcome notation rules out anticipation effects where $D_{g,2}$ could affect the period-one outcome, and dynamic effects where $D_{g,1}$ could affect the period-two outcome. That first difference regressions rule out anticipation and dynamic effects is an important issue, different from that discussed in this paper. To simplify, let us again assume that the model in levels is linear.
Assumption 4
For $t\in\{1,2\}$,
$Y_{t}(d)=Y_{t}(0)+S_{t}d.$ (4)
The slope $S_{t}$ is indexed by $t$, as the treatment’s effect might be time-varying.
Restating Assumption 3, under the linear causal model in levels.
Remember that $\Delta Y_{g}(0)$ is $g$’s potential first-differenced outcome, if $\Delta D_{g}=0\Leftrightarrow D_{g,2}=D_{g,1}$. Then, under the linear causal model in levels in Assumption 4,
$\Delta Y(0)=Y_{2}(D_{1})-Y_{1}(D_{1})=Y_{2}(0)-Y_{1}(0)+(S_{2}-S_{1})D_{1}:$
the residual of the first-differenced model is a function of $D_{1}$. Moreover,
$\Delta Y(\Delta d)=Y_{2}(D_{1}+\Delta d)-Y_{1}(D_{1})=Y_{2}(D_{1}+\Delta d)-Y_{2}(D_{1})+Y_{2}(D_{1})-Y_{1}(D_{1})=\Delta Y(0)+S_{2}\Delta d.$
Thus, Assumption 3 is equivalent to
$cov(\Delta D,Y_{2}(0)-Y_{1}(0)+(S_{2}-S_{1})D_{1})=cov(\Delta D,S_{2})=cov\left(\left(\Delta D\right)^{2},S_{2}\right)=0.$ (5)
A test of a sufficient condition for Assumption 3, under a linear causal model in levels.
In view of (5), a sufficient, more interpretable condition for Assumption 3 is
$cov(\Delta D,Y_{2}(0)-Y_{1}(0))=cov(\Delta D,S_{2})=cov\left(\left(\Delta D\right)^{2},S_{2}\right)=0,$
$\Delta D\perp\!\!\!\perp(S_{2}-S_{1},D_{1}).$ (6)
As $\Delta D$ and $D_{1}$ are observed, (2.3) for instance has the following testable implication:
$cov(\Delta D,D_{1})=0.$ (7)
A simple way to test (7) is to run a linear regression of $\Delta D$ on $D_{1}$. As (7) is a testable implication of a sufficient condition for Assumption 3, Assumption 3 may still hold when (7) is rejected, but that would require that $\Delta D$ be correlated to $S_{2}-S_{1}$ in a way that exactly offsets its correlation with $D_{1}$. For instance, if $E(S_{2}-S_{1}|D_{1},D_{2})=E(S_{2}-S_{1}),$ then
$cov(\Delta D,(S_{2}-S_{1})D_{1})=E(S_{2}-S_{1})cov(\Delta D,D_{1}),$
so assuming that $S_{2}-S_{1}$ is mean independent of $(D_{1},D_{2})$ is not enough to ensure that $cov(\Delta D,(S_{2}-S_{1})D_{1})=0$, if $E(S_{2}-S_{1})\neq 0$ and $cov(\Delta D,D_{1})\neq 0$. Thus, when (7) is rejected, advocating for Assumption 3 requires advocating for exactly offsetting correlations between $\Delta D$ and $S_{2}-S_{1}$ and between $\Delta D$ and $D_{1}$.
If one is ready to assume constant effects over time, Assumption 3 no longer implies strong restrictions.
In view of (5), if one is ready to assume that $S_{2}=S_{1}$, Assumption 3 is equivalent to
$cov(\Delta D,Y_{2}(0)-Y_{1}(0))=cov(\Delta D,S_{2})=cov\left(\left(\Delta D\right)^{2},S_{2}\right)=0,$
an untestable condition very close in spirit to the idea that $\Delta D$ is as good as randomly assigned. Therefore, it is only if treatment effects are time-varying that Assumption 3 implies strong restrictions when combined with a causal model in levels.
2.4 Under a causal model in levels, if $\Delta D$ is not independent of $D_{1}$ and treatment effects change over time, $\beta$ may not identify a convex combination of effects.
In this section, I consider the following assumption:
Assumption 5
$cov(\Delta D,Y_{2}(0)-Y_{1}(0))=cov(\Delta D,S_{2})=cov\left(\left(\Delta D\right)^{2},S_{2}\right)=0,$ (8)
and
$(D_{1},D_{2})\perp\!\!\!\perp\Delta S.$ (9)
Assumption 5 is almost equivalent to (2.3), except that the testable and potentially problematic part of (2.3), involving the independence between $\Delta D$ and $D_{1}$, has been taken out, and replaced by (9), which requires that units’ treatment path $(D_{1},D_{2})$ be independent of the change in their treatment effect $\Delta S$. Thus, Assumption 5 is somewhere in between assuming that the treatment change $\Delta D$ is as good as randomly assigned, and assuming that the treatment path $(D_{1},D_{2})$ is as good as randomly assigned, which is stronger.
Theorem 2
If Assumptions 4 and 5 hold,
$\beta=\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\Delta D)+cov(\Delta D,\Delta SD_{1})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}$
$=$ $\frac{V(\Delta D)E(S_{2})+cov(\Delta D,D_{1})E(\Delta S)}{\sum_{t^{\prime}=1}^{2}V(D_{t^{\prime}})-cov(D_{1},D_{2})}$
$=$ $\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
The first equality follows from Assumption 4. The second follows from Assumption 5. QED.
Theorem 2 shows that under a linear causal model in levels and Assumption 5, $\beta$ estimates a weighted sum of the average slopes $E(S_{1})$ and $E(S_{2})$. Weights are guaranteed to be positive if the treatment is binary. [^2] To see this, note that $V(D_{1})-cov(D_{1},D_{2})=E((D_{1}-D_{2})(D_{1}-E(D_{1})))$, and $(D_{1}-D_{2})(D_{1}-E(D_{1}))\geq 0$ almost surely for binary variables. This is consistent with the results of Athey and Imbens (2022) and Arkhangelsky et al. (2021), who have shown that with a binary randomized treatment, OLS TWFE regressions always estimate a convex combination of effects. By Cauchy-Schwarz’s inequality, weights are also guaranteed to be positive if $V(D_{1})=V(D_{2})$. Finally, weights are also guaranteed to be positive if $cov(D_{1},D_{2})=0$, meaning that the treatments are not serially correlated, or if $cov(D_{1},D_{2})\leq 0$, meaning that the treatments are negatively correlated. On the other hand, one of the two weights could be negative if $D_{1}$ and $D_{2}$ are non-binary, positively correlated, and have different variances.
Instead of Assumption 5, one could assume that
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})$ (10)
or the even stronger condition
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{1}(0),Y_{2}(0),S_{1},S_{2}).$ (11)
Both conditions are sufficient for Assumption 5 to hold. Under (11), estimating a first-difference regression is not necessary anymore: to estimate the treatment’s effect, one can merely regress $Y_{1}$ on an intercept and $D_{1}$ to estimate $E(S_{1})$, and one can regress $Y_{2}$ on an intercept and $D_{2}$ to estimate $E(S_{2})$. Yet, replacing Assumption 5 by (10) or (11) does not lead to a different decomposition from that in Theorem 2: even with a randomly assigned treatment path, and under assumptions under which first-differencing is not necessary anymore, a first-difference regression may not estimate a convex combination of effects under a causal model in levels with time-varying effects. This is simply due to the fact that with a causal model in levels with time-varying effects, the first-difference regression is misspecified. Under (10),
$E(\Delta Y|D_{1},D_{2})=E(Y_{2}(0)-Y_{1}(0))+\Delta D\times E(S_{2})+D_{1}(E(S_{2})-E(S_{1})).$ (12)
In view of (12), correcting for this misspecification is straightforward under (10): one just needs to regress $\Delta Y$ on an intercept, $\Delta D$, and $D_{1}$. The coefficient on $\Delta D$ identifies $E(S_{2})$, while that on $D_{1}$ identifies $E(S_{2})-E(S_{1})$.
3 Application
Data.
I use the industry-level panel data set constructed by Acemoglu et al. (2016). In their Table 2 Column (2), the authors estimate a first-difference regression of industries’ employment evolutions on changes in industries’ Chinese-import-penetration ratio. The regression leverages three time periods and two first differences. To fit in the two-period setting considered in this paper, I estimate a similar regression, with the same weights, clustering standard errors at the same level as the authors, but keeping only two time periods and one first difference, from 1999 to 2007. Doing so, I find that $\widehat{\beta}=-0.78$ (s.e.=$0.22$), which is very close to the coefficient in the paper’s Table 2 Column (2) ($-0.81$, s.e.=$0.16$). According to this regression, an increase in the Chinese-import-penetration ratio decreases US employment.
Test of (7).
I start by testing (7), by regressing $\Delta D$, the change in industries’ import-penetration ratio, on $D_{1}$, industries’ import-penetration ratio in 1999. The coefficient on $D_{1}$ is equal to $0.74$, and it is highly significant (s.e.=$0.16$). Thus, the industries that experienced the largest growth of their imports-penetration ratio from 1999 to 2007 already had a larger imports-penetration ratio in 1999.
Estimating the weights attached to $\beta$ in Theorem 2.
Next, I estimate the weights attached to $\beta$ in Theorem 2, namely under a causal model in levels with time varying effects, and without assuming that $\Delta D$ is independent of $D_{1}$. Doing so, I find that $\beta$ estimates a weighted sum of the average effects of import penetration in 1999 and 2007, where the 2007 average effect receives a weight equal to 1.3, while the 1999 average effect receives a weight equal to -0.3. Thus, under a causal model in levels and Assumption 5, the first-difference regression does not estimate a convex combination of effects, if the effect of import penetration on US employment changes from 1999 to 2007.
Regressing $\Delta Y$ on $\Delta D$ and $D_{1}$.
Under (10), one can estimate the 2007 average effect of import penetration, as well as the difference between the 2007 and 1999 average effect, by regressing $\Delta Y$ on $\Delta D$ and $D_{1}$. The coefficient on $\Delta D$ is equal to $-0.64$ (s.e.=0.24), which is 18% smaller than $\widehat{\beta}$. The coefficient on $D_{1}$ is equal to $-0.45$ and it is marginally significant (s.e. =$0.25$), thus suggesting that the effect of import penetration might have become more negative between 1999 and 2007. Note however that the regression of $\Delta Y$ on $\Delta D$ and $D_{1}$ relies on (10). This condition requires that industries’ import-penetration ratios be randomly assigned in levels, which is stronger than assuming that they are randomly assigned in first differences.