### v4

Abstract
I consider treatment-effect estimation with a two-periods panel, using a first-difference regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. To justify this regression, one may assume that $\Delta D_{g}$ is as good as randomly assigned, namely uncorrelated to the residual of the first-differenced model and to the treatment’s effect. This paper shows that if one posits a causal model in levels between the treatment and the outcome, then the residual of the first-differenced model is a function of the period-one treatment $D_{g,1}$, so the first-difference regression may suffer from an omitted variable bias whenever $\Delta D_{g}$ is correlated to $D_{g,1}$. To solve this problem, one can control for $E(\Delta D_{g}|D_{g,1})$ in the regression. I apply these results to a first-difference regression of the 1999 to 2007 employment evolutions of US industries on the evolution of their Chinese imports penetration ratio, estimated on the data of Acemoglu et al. (2016). The first-difference regression indicates a large negative effect of Chinese imports on US employment, but $\Delta D_{g}$ and $D_{g,1}$ are very strongly correlated, so this regression may suffer from an omitted variable bias. If one controls for $E(\Delta D_{g}|D_{g,1})$, the effect of Chinese imports penetration on US employment is 33% to 41% smaller than in the first-difference regression, and less significant.
1 Introduction
To estimate a treatment’s effect with a two-periods panel data set, researchers often use a first-difference regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. Let $\widehat{\beta}$ denote the coefficient on $\Delta D_{g}$. One may invoke a model-based parallel trends assumption to justify $\widehat{\beta}$. However, under parallel trends, $\widehat{\beta}$ may not estimate a convex combination of treatment effects, outside of the canonical design where all units are untreated at period one and treatment is binary (de Chaisemartin and D’Haultfœuille, 2020; Callaway et al., 2021). An alternative, design-based justification of $\widehat{\beta}$ assumes that $\Delta D_{g}$ is as good as randomly assigned. This paper considers this second justification.
Assume that one posits a causal model in first difference, where different values of the treatment evolution $\Delta D_{g}$ may give rise to different counterfactual values of the outcome evolution $\Delta Y_{g}(\Delta d)$. Then, it is well-known that if $\Delta D_{g}$ is as-good-as randomly assigned, namely uncorrelated to the residual of the first-differenced model and to the effect of the first-differenced treatment, $\widehat{\beta}$ estimates an average effect of the first-differenced treatment on the first-differenced outcome. However, the first-difference specification does not always come from the researcher’s belief that the causal relation between the treatment and the outcome only arises in first difference. Rather, it may sometimes come from the researcher’s belief that first-differencing is necessary to net out time-invariant counfounders.
If one instead posits a causal model in levels, where different values of the treatment $D_{g,t}$ may give rise to different counterfactual values of the outcome $Y_{g,t}(d)$, I show that the residual of the first-differenced model is a function of $D_{g,1}$, so $\Delta D_{g}$ uncorrelated to that residual essentially implies that $\Delta D_{g}$ is uncorrelated to $D_{g,1}$. This is a strong condition, which can be tested by regressing $\Delta D_{g}$ on $D_{g,1}$. If $\Delta D_{g}$ is correlated to $D_{g,1}$, $\widehat{\beta}$ suffers from an omitted variable bias if the treatment’s effect changes over time. Intuitively, due to the correlation between $\Delta D_{g}$ and $D_{g,1}$, $\widehat{\beta}$ attributes to $\Delta D_{g}$ changes in the outcome that would have happened without any treatment change, and that are caused by changes in the treatment’s effect. Due to this omitted variable bias, I show that $\widehat{\beta}$ may estimate a non-convex combination of effects, even if the treatment paths $(D_{g,1},D_{g,2})$ are randomly assigned, which is stronger than assuming that $\Delta D_{g}$ is randomly assigned.
Then, I show that these issues can be fixed by regressing $\Delta Y_{g}$ on $\Delta D_{g}$, controlling for $E(\Delta D_{g}|D_{g,1})$. If $\Delta D_{g}$ is randomly assigned conditional on $D_{g,1}$, the coefficient on $\Delta D_{g}$, $\beta_{D_{1}}$, identifies a variance-weighted average of conditional averages of the treatment effect at period two (Angrist, 1998). For estimation, I propose a classical two-step estimator, where in a first step one estimates $E(\Delta D_{g}|D_{g,1})$ non parametrically. Alternatively, one could prefer using a doubly-robust non-parametric estimator, where the tuning parameters are chosen automatically, in a data-driven manner. However, doubly-robust estimation of $\beta_{D_{1}}$ has to rely on some assumptions: there needs to be stayers, namely observations such that $\Delta D_{g}=0$, and one has to assume that the treatment effect is mean-independent of $D_{g,1}$. Then, one can show that $\beta_{D_{1}}$ solves a moment condition that is doubly robust in the non-parametric nuisance functions $E(\Delta D_{g}|D_{g,1})$ and $E(\Delta Y_{g}|\Delta D_{g}=0,D_{g,1})$ (Robinson, 1988), and one can follow, say, Chernozhukov et al. (2018) to propose a doubly-robust estimator of $\beta_{D_{1}}$ with data-driven choices of the tuning parameters. However, this estimator requires stayers, and it restricts treatment-effect heterogeneity. An alternative would be to consider the Augmented Inverse Propensity Weighted (AIPW) estimand of Robins et al. (1994), which can be estimated in a doubly-robust manner, without making any assumption on the treatment’s effect. However, this estimand requires that $\Delta D$ be binary, while $\Delta D$ is often continuous in applications where first-difference regressions are used. Then, the generalization of the AIPW estimand to continuous treatments considered by Colangelo and Lee (2020) may be an appealing option. However, that paper proposes an estimator of the dose-response function, which converges at a slow, non-parametric rate, but does not propose a method to aggregate the dose-response function into a scalar that can be estimated at the parametric rate.
As mentioned earlier, the issues highlighted in this paper can be avoided by assuming that there does not exist a causal model in levels between the treatment and the outcome. Then, researchers should make this assumption explicit, and should defend why the theory underlying their estimation delivers a causal model in first difference but not in levels.
As an empirical application, I revisit Acemoglu et al. (2016). One of their regressions is a first-difference regression with three periods and two first-differences, to estimate the effect of US industries’ Chinese-imports-penetration ratio on their employment. To fit in the setting considered in this paper, I reestimate their regression, keeping only two time periods, 1999 and 2007, and one first difference. Doing so, I obtain a large and negative coefficient $\widehat{\beta}$, very close to that in the paper, suggesting a negative effect of Chinese imports on US employment. I start by regressing $\Delta D_{g}$ on $D_{g,1}$, and I find that the two variables are strongly positively correlated: industries where the import-penetration ratio increased the most already had a larger import-penetration ratio in 1999. Then, $\widehat{\beta}$ could suffer from an omitted variable bias. Indeed, I show that even if the paths of Chinese-imports-penetration ratios $(D_{g,1},D_{g,2})$ are randomly assigned, $\widehat{\beta}$ estimates a non-convex combination of the treatment effects in 1999 and 2007 where the 1999 effect receives a weight equal to $-0.30$. Finally, I turn to my alternative estimator $\widehat{\beta}_{D_{1}}$. This application does not have stayers, so I use a two-step estimator where $E(\Delta D_{g}|D_{g,1})$ is estimated non-parametrically in the first step, using a local linear kernel regression. $\widehat{\beta}_{D_{1}}$ is 41% smaller than $\widehat{\beta}$, it is marginally insignificant (t-stat=-1.59), and it is significantly different from $\widehat{\beta}$ at the 10% level. Using instead a series estimator to estimate $E(\Delta D_{g}|D_{g,1})$, $\widehat{\beta}_{D_{1}}$ is 33% smaller than $\widehat{\beta}$, and it is marginally significant (t-stat=-1.70).
Some of the ideas in this paper originated while working on de Chaisemartin and Lei (2021), see in particular Sections 2.1 and 3.3 therein. That paper considers first-difference 2SLS Bartik regressions, that are sometimes justified by assuming that the first-difference shocks used to construct the Bartik instrument are as good as randomly assigned (Borusyak et al., 2022). As noted in de Chaisemartin and Lei (2021), issues with design-based justifications of first-difference regressions are not specific to 2SLS Bartik regressions, which is why these ideas are now presented in this standalone paper. Future versions of de Chaisemartin and Lei (2021) will not contain material overlapping with that in this paper. Another related paper is Klosin (2024), who shows that fixed effects regressions, which are closely related to first-difference ones, can yield inconsistent treatment-effect estimators, even with randomly assigned treatments, when past outcomes have an effect on the current outcome. In the setting I consider, the first-difference regression can be misleading even if past outcomes do not have an effect on the current outcome.
2 Set-up
Two-periods panel.
One is interested in estimating the effect of a treatment on an outcome, using a panel data set, with $G$ independent and identically distributed units indexed by $g$, and with two time periods. For $t\in\{1,2\}$, let $Y_{g,t}$ and $D_{g,t}$ respectively denote the outcome and treatment of $g$ at $t$.
Assumption 1
$(Y_{g,1},Y_{g,2},D_{g,1},D_{g,2})_{g=1,...,G}$ are i.i.d.
As units are assumed i.i.d., I sometimes drop the $g$ subscript below.
First-difference regression.
Let $\Delta$ denote the first difference operator. To estimate the treatment’s effect, researchers often regress $\Delta Y_{g}$ on an intercept and $\Delta D_{g}$, leading to
$\widehat{\beta}:=\frac{\sum_{g}(\Delta D_{g}-\Delta D_{.})\Delta Y_{g}}{\sum_{g}(\Delta D_{g}-\Delta D_{.})^{2}},$
where $\Delta D_{.}=1/G\sum_{g}\Delta D_{g}.$ Assuming that $(\Delta Y,\Delta D)$ has a finite second moment, under Assumption 1 the probability limit of $\widehat{\beta}$ when $G\rightarrow+\infty$ is equal to
$\beta:=\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}.$
3 Potential problem
3.1 Under a linear causal model in first-difference, $\beta$ identifies an average slope if the first-differenced treatment is as good as randomly assigned.
A linear causal model in first difference.
In this section, I consider a causal model in first-difference, where different values of the treatment evolution can lead to different counterfactual outcome evolutions. Accordingly, let $\Delta Y_{g}(\Delta d)$ denote $g$’s potential first-differenced outcome if $\Delta D_{g}=\Delta d$. $\Delta Y_{g}(0)$ is $g$’s potential first-differenced outcome if $\Delta D_{g}=0\Leftrightarrow D_{g,2}=D_{g,1}$, meaning that $g$’s treatment does not change. I assume that the treatment evolution has a linear effect on the outcome evolution.
Assumption 2
$\Delta Y(\Delta d)=\Delta Y(0)+S_{\delta}\times\Delta d.$ (1)
Assumption 2 is a linear effect assumption, but it allows for heterogeneous effects across units: the slope of the potential first-differenced outcome function $S_{\delta}$ is a random variable, that may vary across units. Under Assumption 2,
$\Delta Y=\Delta Y(0)+S_{\delta}\times\Delta D.$ (2)
As good as randomly assigned first-differenced treatment.
Assumption 3
$cov(\Delta D,\Delta Y(0))=cov(\Delta D,S_{\delta})=cov\left(\left(\Delta D\right)^{2},S_{\delta}\right)=0.$ (3)
Assumption 3 requires that the first-differenced treatment be uncorrelated to units’ outcome evolutions without a treatment change, the residual in the model relating the first-differenced outcome and the first-differenced treatment. Assumption 3 also requires that the first-differenced treatment and its square be uncorrelated to the slope $S_{\delta}$.
Theorem 1
If Assumptions 2 and 3 hold, $\beta=E(S_{\delta})$.
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,\Delta Y(0))+cov(\Delta D,S_{\delta}\times\Delta D)}{V(\Delta D)}$
$=$ $\frac{E\left(S_{\delta}\times\left(\Delta D\right)^{2}\right)-E(\Delta D)E(S_{\delta}\times\Delta D)}{V(\Delta D)}$
$=$ $E(S_{\delta}).$
The first equality follows from (2). The second and third equalities follow from Assumption 3. QED.
Theorem 1 is a well-known result. It shows that $\beta$ identifies the average slope $E(S_{\delta})$ under Assumptions 2 and 3. Under Assumption 3, $\beta$ can still identify a well-defined causal effect if the effect is non-linear: the linearity assumption is made to simplify the discussion.
3.2 Unpacking Assumption 3 under a causal model in levels.
A linear causal model in levels.
Assume that instead of positing that different counterfactual treatment evolutions can lead to different outcome evolutions, one prefers to posit that different counterfactual treatment levels can lead to different outcome levels. Then, let $Y_{g,t}(d)$ denote $g$’s potential outcome if $D_{g,t}=d$. $Y_{g,t}(0)$ is $g$’s potential outcome at $t$ if $D_{g,t}=0$, namely without treatment. [^1] Implicitly, our potential outcome notation rules out anticipation effects where $D_{g,2}$ could affect the period-one outcome, and dynamic effects where $D_{g,1}$ could affect the period-two outcome. That first difference regressions rule out anticipation and dynamic effects is an important issue, different from that discussed in this paper. To simplify, let me again assume that the model in levels is linear.
Assumption 4
For $t\in\{1,2\}$,
$Y_{t}(d)=Y_{t}(0)+S_{t}d.$ (4)
The slope $S_{t}$ is indexed by $t$, as the treatment’s effect might be time-varying. Under (4), one has
$\Delta Y=Y_{2}(0)+S_{2}D_{2}-Y_{1}(0)-S_{1}D_{1}=Y_{2}(0)-Y_{1}(0)+S_{2}\times\Delta D+\Delta S\times D_{1},$ (5)
a decomposition I use repeatedly in what follows.
Restating Assumption 3, under the linear causal model in levels.
Remember that $\Delta Y_{g}(0)$ is $g$’s potential first-differenced outcome, if $\Delta D_{g}=0\Leftrightarrow D_{g,2}=D_{g,1}$. Then, under the linear causal model in levels in Assumption 4,
$\Delta Y(0)=Y_{2}(D_{1})-Y_{1}(D_{1})=Y_{2}(0)-Y_{1}(0)+\Delta S\times D_{1}:$
the residual of the first-differenced model is a function of $D_{1}$. Moreover,
$\Delta Y(\Delta d)=Y_{2}(D_{1}+\Delta d)-Y_{1}(D_{1})=Y_{2}(D_{1}+\Delta d)-Y_{2}(D_{1})+Y_{2}(D_{1})-Y_{1}(D_{1})=\Delta Y(0)+S_{2}\Delta d.$
Thus, Assumption 3 is equivalent to
$cov(\Delta D,Y_{2}(0)-Y_{1}(0)+\Delta S\times D_{1})=cov(\Delta D,S_{2})=cov\left(\left(\Delta D\right)^{2},S_{2}\right)=0.$ (6)
A test of a sufficient condition for Assumption 3, under a linear causal model in levels.
In view of (6), a sufficient, more interpretable condition for Assumption 3 is
$cov(\Delta D,Y_{2}(0)-Y_{1}(0))=cov(\Delta D,S_{2})=cov\left(\left(\Delta D\right)^{2},S_{2}\right)=0,$ (7)
$\Delta D\perp\!\!\!\perp(\Delta S,D_{1}).$ (8)
As $\Delta D$ and $D_{1}$ are observed, (8) for instance has the following testable implication:
$cov(\Delta D,D_{1})=0.$ (9)
A simple way to test (9) is to run a linear regression of $\Delta D$ on $D_{1}$. As (9) is a testable implication of a sufficient condition for Assumption 3, Assumption 3 may still hold when (9) is rejected, but that would require that $\Delta D$ be correlated to $\Delta S$ in a way that exactly offsets its correlation with $D_{1}$. For instance, if
$E(\Delta S|D_{1},D_{2})=E(\Delta S),$ (10)
then
$cov(\Delta D,\Delta S\times D_{1})=E(\Delta S)cov(\Delta D,D_{1}),$ (11)
so assuming that $\Delta S$ is mean independent of $(D_{1},D_{2})$ is not enough to ensure that $cov(\Delta D,\Delta S\times D_{1})=0$, if $E(\Delta S)\neq 0$ and $cov(\Delta D,D_{1})\neq 0$. Thus, when (9) is rejected, advocating for Assumption 3 requires advocating for exactly offsetting correlations between $\Delta D$ and $\Delta S$ and between $\Delta D$ and $D_{1}$.
3.3 Omitted variable bias when $cov(\Delta D,D_{1})\neq 0$.
OVB formula.
If (7) holds but (8) does not, then it is easy to see that
$\beta=E(S_{2})+\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)},$
a standard omitted variable bias (OVB) formula. If one is ready to assume that the treatment effect does not change over time ($S_{2}=S_{1}$), then the OVB is equal to zero and $\beta=E(S_{2})$. But if the treatment effect changes over time, the OVB may differ from zero. If $E(\Delta S|D_{1},D_{2})=E(\Delta S),$ (11) shows that the OVB is positive if $E(\Delta S)$ and $cov(\Delta D,D_{1})$ are of the same sign, and negative otherwise. For instance, if $cov(\Delta D,D_{1})>0$ and the treatment effect increases over time, $\beta$ is larger than $E(S_{2})$. Intuitively, due to the correlation between $\Delta D$ and $D_{1}$, $\beta$ mistakenly attributes to $\Delta D$ increases in the outcome that would have happened even without any treatment change, due to increases in the treatment’s effect.
$\beta$ may identify a non-convex combination of $E(S_{1})$ and $E(S_{2})$.
As the OVB in the previous display is a function of $\Delta S$, one may wonder if, at the end of the day, $\beta$ still identifies a relatively interpretable causal effect, like a convex combination of the treatment effects at period one and two. The following result shows this is not necessarily the case.
Theorem 2
If Assumption 4, (7), and (10) hold,
$\beta=\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\times\Delta D)+cov(\Delta D,\Delta S\times D_{1})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}$
$=$ $\frac{V(\Delta D)E(S_{2})+cov(\Delta D,D_{1})E(\Delta S)}{\sum_{t^{\prime}=1}^{2}V(D_{t^{\prime}})-cov(D_{1},D_{2})}$
$=$ $\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
The first equality follows from (5). The second follows from (7) and (10). QED.
Theorem 2 shows that under a linear causal model in levels, (7), and (10), $\beta$ estimates a weighted sum of the average slopes $E(S_{1})$ and $E(S_{2})$. Weights are guaranteed to be positive if the treatment is binary. [^2] To see this, note that $V(D_{1})-cov(D_{1},D_{2})=E((D_{1}-D_{2})(D_{1}-E(D_{1})))$, and $(D_{1}-D_{2})(D_{1}-E(D_{1}))\geq 0$ almost surely for binary variables. This is consistent with the results of Athey and Imbens (2022) and Arkhangelsky et al. (2021), who have shown that with a binary randomized treatment, OLS TWFE regressions always estimate a convex combination of effects. By Cauchy-Schwarz’s inequality, weights are also guaranteed to be positive if $V(D_{1})=V(D_{2})$. Finally, weights are also guaranteed to be positive if the treatments are not serially correlated ($cov(D_{1},D_{2})=0$) or negatively correlated ($cov(D_{1},D_{2})<0$). On the other hand, one of the two weights could be negative if $D_{1}$ and $D_{2}$ are non-binary, positively correlated, and have different variances. Instead of (7) and (10), one could assume that
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})$ (12)
or the even stronger condition
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{1}(0),Y_{2}(0),S_{1},S_{2}).$ (13)
Both conditions are sufficient for (7) and (10) to hold. Under (13), estimating a first-difference regression is not necessary anymore: to estimate the treatment’s effect, one can merely regress $Y_{1}$ on an intercept and $D_{1}$ to estimate $E(S_{1})$, and one can regress $Y_{2}$ on an intercept and $D_{2}$ to estimate $E(S_{2})$. Yet, replacing (7) and (10) by (12) or (13) does not lead to a different decomposition from that in Theorem 2: even with a randomly assigned treatment path, and under assumptions under which first-differencing is not necessary anymore, a first-difference regression may not estimate a convex combination of effects.
4 Solutions
4.1 Identification
Let $\beta_{D_{1}}$ be the coefficient on $\Delta D$ in a regression of $\Delta Y$ on an intercept, $\Delta D$, and $E(\Delta D|D_{1})$.
Identifying assumption.
I consider the following assumption.
Assumption 5
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})|D_{1},$ (14)
Assumption 5 requires random assignment of $\Delta D$ conditional on $D_{1}$. Let $\Delta D_{r}=\Delta D-E(\Delta D|D_{1})$ denote the residual from the CEF decomposition of $\Delta D$ with respect to $D_{1}$.
Theorem 3
If Assumptions 4 and 5 hold,
$\beta_{D_{1}}=E\left(\frac{V(\Delta D|D_{1})}{E\left(V(\Delta D|D_{1})\right)}E(S_{2}|D_{1})\right).$
Proof
Under Assumption 5,
$cov(\Delta D_{r},Y_{2}(0)-Y_{1}(0))=cov(\Delta D_{r},S_{2})=cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))$
$=$ $cov\left(\Delta D_{r},\Delta S\times D_{1}\right)=0.$ (15)
For instance,
$cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))=$ $E(\Delta D_{r}S_{2}E(\Delta D|D_{1}))$
$=$ $E(E(\Delta D_{r}|S_{2},D_{1})S_{2}E(\Delta D|D_{1}))$
$=$ $E((E(\Delta D|S_{2},D_{1})-E(\Delta D|D_{1}))S_{2}E(\Delta D|D_{1}))$
$=$ $0$
Then,
$\beta_{D_{1}}=$ $\frac{cov(\Delta D_{r},\Delta Y)}{V(\Delta D_{r})}$
$=$ $\frac{cov(\Delta D_{r},Y_{2}(0)-Y_{1}(0))+cov(\Delta D_{r},S_{2}\times\Delta D)+cov(\Delta D_{r},\Delta S\times D_{1})}{V(\Delta D_{r})}$
$=$ $\frac{cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))+cov(\Delta D_{r},S_{2}\times\Delta D_{r})}{V(\Delta D_{r})}$
$=$ $\frac{E\left((\Delta D_{r})^{2}S_{2}\right)}{E\left((\Delta D_{r})^{2}\right)}$
$=$ $E\left(\frac{V(\Delta D|D_{1})}{E\left(V(\Delta D|D_{1})\right)}E(S_{2}|D_{1})\right).$
The first equality follows from the Frich-Waugh-Lovell theorem. The second follows from (5). The third and fourth follow from (4.1). The last follows from the law of iterated expectations and Assumption 5.
QED.
The result says that under Assumption 5, $\beta_{D_{1}}$ identifies a weighted average of conditional average slopes $E(S_{2}|D_{1})$. This is essentially a restatement of the well-known result in Angrist (1998) in the setting I consider. [^3] In this section, I still make a linear treatment effect assumption, as in the previous section. This simplifies the analysis and makes it easier to compare the results therein to those in the previous section. Of course, while the linearity assumption was not problematic to derive the “negative” results in the previous section, it becomes more problematic for the more “positive” results in this section. The result in Angrist (1998) has been generalized to a non-binary treatment without making a linearity assumption. Thus, the linear treatment effect model is not of essence for Theorem 3.
4.2 Estimation
When $D_{1}$ is a discrete variable, estimation of $E(\Delta D|D_{1})$ and then of $\beta_{D_{1}}$ is straightforward. Accordingly, in this section I focus on the case where $D_{1}$ is continuous.
4.2.1 Classical two-step estimator
To estimate $\beta_{D_{1}}$, one can start by estimating $E(\Delta D|D_{1})$. Then, one lets $\widehat{\Delta D}_{r}=\Delta D-\widehat{E}(\Delta D|D_{1})$, and one regresses $\Delta Y$ on $\widehat{\Delta D}_{r}$. $E(\Delta D|D_{1})$ may be estimated non-parametrically, say using local-linear kernel regressions. The properties of two-steps estimators with a non-parametric first-step estimator are well understood (Newey, 1994). Alternatively, $E(\Delta D|D_{1})$ may be estimated parametrically, assuming a parametric functional-form for $E(\Delta D|D_{1})$, say a polynomial of a certain order. Then researchers can assess results’ sensitivity to the chosen functional form, by varying the polynomial order.
4.2.2 Doubly-robust estimation with stayers, under a restriction on treatment-effect heterogeneity
An issue with classical two-step estimators is that tuning parameters (the bandwidths in the local-linear kernel regressions, or the polynomial order) are not chosen in a data-driven manner. On that front, estimators solving moment conditions that are doubly-robust in the non-parametric nuisance functions offer an appealing alternative. However, doubly-robust estimation of $\beta_{D_{1}}$ requires that there exist stayers, namely units such that $\Delta D=0$, and it also requires making a strong assumption on the treatment’s effect.
Let stayers be units such that $\Delta D=0$. In this section, I assume that $P(\Delta D=0|D_{1})$, the proportion of stayers conditional on $D_{1}$, is bounded by below by a strictly positive $\epsilon$.
Assumption 6
$P(\Delta D=0|D_{1})\geq\varepsilon>0$.
Using standard properties of residuals from CEF decompositions,
$\beta_{D_{1}}=\frac{E((\Delta D-E(\Delta D|D_{1}))\Delta Y)}{E((\Delta D-E(\Delta D|D_{1}))\Delta D)}=\frac{E((\Delta D-E(\Delta D|D_{1}))(\Delta Y-E(\Delta Y|D_{1},\Delta D=0)))}{E((\Delta D-E(\Delta D|D_{1}))\Delta D)}.$
Therefore, letting $f_{0}(D_{1})=E(\Delta D|D_{1})$ and $g_{0}(D_{1})=E(\Delta Y|D_{1},\Delta D=0)$, $\beta_{D_{1}}$ solves
$E((\Delta D-f_{0}(D_{1}))(\Delta Y-g_{0}(D_{1})-\beta_{D_{1}}\Delta D))=0.$ (16)
Under Assumptions 4 and 5 alone, this moment condition is not doubly-robust in $(f_{0}(D_{1}),g_{0}(D_{1}))$: it holds at $(f_{0}(D_{1}),g(D_{1}))$ for any $g$, but it does not hold at $(f(D_{1}),g_{0}(D_{1}))$ for any $f$. However, it is doubly-robust under the following assumption.
Assumption 7
$E(S_{2}|D_{1})=E(S_{2})$.
Assumption 7 requires that $S_{2}$ be mean independent of $D_{1}$. This is more in the spirit of assuming random-assignment in levels than in first-difference.
I now show that (16) is a doubly-robust moment condition under Assumptions 4, 5, and 7. First, it is easy to see that for all $g(.)$,
$E((\Delta D-f_{0}(D_{1}))(\Delta Y-g(D_{1})-\beta_{D_{1}}\Delta D))=0.$
Second, under (5) and Assumption 5,
$g_{0}(D_{1})=E(Y_{2}(0)-Y_{1}(0)|D_{1})+D_{1}E(\Delta S|D_{1}).$
Therefore,
$\Delta Y-g_{0}(D_{1})-\beta_{D_{1}}\Delta D$
$=$ $Y_{2}(0)-Y_{1}(0)-E(Y_{2}(0)-Y_{1}(0)|D_{1})+D_{1}(\Delta S-E(\Delta S|D_{1}))+\Delta D(S_{2}-\beta_{D_{1}}).$
Under Assumption 7, $\beta_{D_{1}}=E(S_{2})$. Therefore,
$E(\Delta Y-g_{0}(D_{1})-\beta_{D_{1}}\Delta D|\Delta D,D_{1})=0$
under Assumptions 5 and 7. Accordingly, for all $f(.)$,
$E((\Delta D-f(D_{1}))(\Delta Y-g_{0}(D_{1})-\beta_{D_{1}}\Delta D))=0.$
Then, one can apply the Double-Debiased machine learning estimator proposed by Chernozhukov et al. (2018) to the moment condition in (16), which is equivalent to the moment condition attached to the partially linear regression model of Robinson (1988). As the non-parametric nuisance functions in the moment condition are CEFs with respect to only one conditioning variable, $D_{1}$, it follows from results in Andrews (1991) that one can just use a polynomial with a cross-validated order to estimate them.
Doubly-robust estimation of $\beta_{D_{1}}$ requires that there exist stayers and also requires making a strong assumption on the treatment’s effect. An alternative would be to consider the Augmented Inverse Propensity Weighted (AIPW) estimand of Robins et al. (1994), which can be estimated in a doubly-robust manner, without making any assumption on the treatment’s effect. However, this estimand requires that $\Delta D$ be binary, while $\Delta D$ is often continuous in applications where first-difference regressions are used. Then, the generalization of the AIPW estimand to continuous treatments considered by Colangelo and Lee (2020) may be an appealing option. However, that paper proposes an estimator of the dose-response function, which converges at a slow, non-parametric rate, but it does not propose a method to aggregate the dose-response function into a scalar that can be estimated at the parametric rate.
4.3 Placebo tests.
Placebo tests are an important step in establishing the credibility of an identifying assumption in observational studies (Imbens et al., 2001; Imbens and Xu, 2024). However, placebo testing Assumption 5 is not straightforward. Assume one observes another period of data, period 0, and let $\Delta Y_{-1}$ and $\Delta D_{-1}$ respectively denote the outcome and treatment evolutions from period 00 to $1$. The coefficient on $\Delta D$ in the regression of $\Delta Y_{-1}$ on an intercept, $\Delta D$, and $E(\Delta D|D_{1})$ could differ from zero, even if Assumption 5 holds with all the unobservables switched back by one period, if $\Delta D$ is correlated to $\Delta D_{-1}$. Instead, I now propose a partly placebo-testable assumption under which a weighted average of conditional slopes is identified. Let $\beta_{D_{0},D_{1}}$ denote the coefficient on $\Delta D$ in a regression of $\Delta Y$ on an intercept, $\Delta D$, and $E(\Delta D|D_{0},D_{1})$. Let $\Delta D_{r,2}=\Delta D-E(\Delta D|D_{0},D_{1})$ denote the residual from the CEF decomposition of $\Delta D$ with respect to $(D_{0},D_{1})$. Let $\beta^{pl}_{D_{0},D_{1}}$ denote the coefficient on $\Delta D$ in a regression of $\Delta Y_{-1}$ on an intercept, $\Delta D$, and $E(\Delta D|D_{0},D_{1})$.
Assumption 8
$\Delta D\perp\!\!\!\perp(Y_{1}(0)-Y_{0}(0),Y_{2}(0)-Y_{1}(0),S_{0},S_{1},S_{2})|D_{0},D_{1}$ (17)
Assumption 8 generalizes Assumption 5 to settings with three dates, by assuming that $\Delta D$ is independent of potential outcome evolutions without treatment and treatment effects at all dates, conditional on the period zero and one treatments.
Theorem 4
If Assumptions 4 and 8 hold,
$\beta_{D_{0},D_{1}}=E\left(\frac{V(\Delta D|D_{0},D_{1})}{E\left(V(\Delta D|D_{0},D_{1})\right)}E(S_{2}|D_{0},D_{1})\right),$
and $\beta^{pl}_{D_{0},D_{1}}=0.$
Proof
The proof of the first equality is similar to the proof of the first equality in Theorem 3, and is therefore omitted. Turning to the second equality,
$\beta^{pl}_{D_{0},D_{1}}=$ $\frac{cov(\Delta D_{r,2},\Delta Y_{-1})}{V(\Delta D_{r,2})}$
$=$ $\frac{cov(\Delta D_{r,2},Y_{1}(0)-Y_{0}(0))+cov(\Delta D_{r,2},S_{1}\Delta D_{-1})+cov(\Delta D_{r,2},\Delta S_{-1}D_{0})}{V(\Delta D_{r,2})}$
$=$ $\frac{cov(\Delta D_{r,2},S_{1}D_{1})-cov(\Delta D_{r,2},S_{1}D_{0})}{V(\Delta D_{r,2})}$
$=$ $0$
QED.
5 Application
Data.
I use the industry-level panel data set constructed by Acemoglu et al. (2016). In their Table 2 Column (2), the authors estimate a first-difference regression of industries’ employment evolutions on changes in industries’ Chinese-import-penetration ratio. The regression leverages three time periods and two first differences. To fit in the two-period setting considered in this paper, I estimate a similar regression, with the same weights, clustering standard errors at the same level as the authors, but keeping only the last two time periods in their dataset, 1999 and 2007, and one first difference. Doing so, I find that $\widehat{\beta}=-0.78$ (s.e.=$0.22$), which is very close to the coefficient in the paper’s Table 2 Column (2) ($-0.81$, s.e.=$0.16$). According to this regression, an increase in the Chinese-import-penetration ratio decreases US employment.
Test of (9).
I start by testing (9), by regressing $\Delta D$, the change in industries’ import-penetration ratio, on $D_{1}$, industries’ import-penetration ratio in 1999. The coefficient on $D_{1}$ is equal to $0.74$, and it is highly significant (s.e.=$0.16$). Thus, the industries that experienced the largest growth of their imports-penetration ratio from 1999 to 2007 already had a larger imports-penetration ratio in 1999.
Estimating the weights attached to $\beta$ in Theorem 2.
Next, I estimate the weights attached to $\beta$ in Theorem 2, namely under a causal model in levels with time varying effects, and without assuming that $\Delta D$ is independent of $D_{1}$. Doing so, I find that $\beta$ estimates a weighted sum of the average effects of import penetration in 1999 and 2007, where the 2007 average effect receives a weight equal to 1.3, while the 1999 average effect receives a weight equal to -0.3. Thus, under a causal model in levels, (7), and (10), the first-difference regression does not estimate a convex combination of effects, if the effect of import penetration on US employment changes from 1999 to 2007.
Estimation of $\beta_{D_{1}}$.
There are only five “stayers” in this application, namely industries whose Chinese-import-penetration ratio does not change over time. Therefore, a doubly-robust estimator cannot be used. Then, I start by letting $\widehat{\Delta D}_{r}=\Delta D-\widehat{E}(\Delta D|D_{1})$, where $\widehat{E}(\Delta D|D_{1})$ is estimated non-parametrically using local linear kernel regressions, with Stata’s npregress kernel command, and the command’s default choices of kernel (epanechnikov) and bandwidth (minimizer of the integrated mean-squared error). Then, I regress $\Delta Y$ on $\widehat{\Delta D}_{r}$. $\widehat{\beta}_{D_{1}}$ is equal to $-0.46$, it is marginally insignificant (s.e. estimated by bootstrap=0.29, t-stat=-1.59), and it is significantly different from $\widehat{\beta}$ at the 10% level (t-stat=-1.89). Using instead a series estimator to estimate $E(\Delta D|D_{1})$, with Stata’s npregress series command, leads to a point estimate of $-0.52$, which is marginally significant (s.e. estimated by bootstrap=0.30, t-stat=-1.70).