### v6

Abstract
We consider treatment-effect estimation using a first-difference regression of an outcome evolution $\Delta Y$ on a treatment evolution $\Delta D$. Under a causal model in levels, the residual of the first-difference regression is a function of the period-one treatment $D_{1}$. Then, researchers should test if $\Delta D$ and $D_{1}$ are correlated: if they are, the first-difference regression may suffer from an omitted variable bias. To solve it, researchers may control for $E(\Delta D|D_{1})$. We apply these results to regressions of US industries’ employment evolutions on the evolution of their Chinese imports, estimated on the data of Acemoglu et al. (2016). $\Delta D$ and $D_{1}$ are strongly correlated. Controlling for $E(\Delta D|D_{1})$ halves the estimated effect of Chinese imports.
1 Introduction
To estimate a treatment’s effect with a two-periods panel, when units’ exposure to treatment varies at period one researchers often use a first-difference (FD) regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. Let $\widehat{\beta}$ denote the coefficient on $\Delta D_{g}$. To justify $\widehat{\beta}$, two identifying assumptions are available in the econometrics literature. First, one may invoke a model-based parallel trends assumption. However, if the treatment effect is heterogeneous, parallel trends is not enough to have that $\widehat{\beta}$ estimates a convex combination of effects whenever one departs from the canonical design where treatment is binary and no one is treated at period one (de Chaisemartin and D’Haultfœuille, 2020). Second, one may assume that $\Delta D_{g}$ is as good as randomly assigned, namely uncorrelated to the effect of $\Delta D_{g}$, and to the residual of a regression of $\Delta Y_{g}$ and $\Delta D_{g}$. This paper considers this second, design-based justification.
Assume that $Y_{g,t}$, the outcome of unit $g$ at $t$, is a linear function of its treatment $D_{g,t}$:
$Y_{g,t}=\alpha_{t}+S_{t}D_{g,t}+u_{g,t}.$ (1)
$S_{t}$ is the effect of increasing the treatment by one unit on the outcome at period $t$, assumed to be constant across units but allowed to vary over time. $\alpha_{t}$ is the expected outcome without any treatment at $t$, and $u_{g,t}$ is a mean-zero residual. First-differencing (1), and letting $v_{g}=D_{g,1}(S_{2}-S_{1})+\Delta u_{g}$,
$\Delta Y_{g}=$ $\alpha_{2}+S_{2}D_{g,2}+u_{g,2}-(\alpha_{1}+S_{1}D_{g,1}+u_{g,1})$
$=$ $\Delta\alpha+S_{2}\Delta D_{g}+D_{g,1}(S_{2}-S_{1})+\Delta u_{g}$
$=$ $\Delta\alpha+S_{2}\Delta D_{g}+v_{g}.$ (2)
Then, under (1), the residual of a regression of $\Delta Y_{g}$ on $\Delta D_{g}$ is $v_{g}=D_{g,1}(S_{2}-S_{1})+\Delta u_{g}$, and $\Delta D_{g}$ has to be uncorrelated to $v_{g}$ for the regression not to suffer from an omitted variable bias (OVB). But if the treatment effect changes over time ($S_{2}\neq S_{1}$), $v_{g}$ is a function of $D_{g,1}$, so having that $\Delta D_{g}$ and $v_{g}$ are uncorrelated essentially requires that $\Delta D_{g}$ and $D_{g,1}$ are uncorrelated. This a strong condition: for instance, if $D_{g,1}$ and $D_{g,2}$ have the same bounded support, $\Delta D_{g}$ and $D_{g,1}$ cannot be independent. This is also a testable condition, and our first recommendation is that researchers invoking random assignment to justify $\widehat{\beta}$ start by regressing $\Delta D_{g}$ on $D_{g,1}$. This is akin to a standard balancing check, where one regresses the treatment on proxies for the outcomes’s residual. But while usually, the set of proxies is open to interpretation, here the residual is a function of $D_{g,1}$, leaving little room for interpretation: $D_{g,1}$ needs to be included in the balancing checks.
Here is some intuition on the OVB that may arise when $\Delta D_{g}$ and $D_{g,1}$ are correlated. Due to the correlation between $\Delta D_{g}$ and $D_{g,1}$, $\widehat{\beta}$ attributes to $\Delta D_{g}$ changes in $Y_{g,t}$ that would have happened without any treatment change, that are caused by changes in the treatment’s effect, the $D_{g,1}(S_{2}-S_{1})$ term in $v_{g}$. Below, we give further intuition in the context of our empirical example. Due to this OVB, we show that $\widehat{\beta}$ may estimate a non-convex combination of effects, even if the treatment paths $(D_{g,1},D_{g,2})$ are randomly assigned, which is stronger than assuming that $\Delta D_{g}$ is randomly assigned.
If $\Delta D_{g}$ and $D_{g,1}$ are correlated and researchers are not willing to assume a constant effect over time, our second recommendation is that they may estimate the treatment’s effect by regressing $\Delta Y_{g}$ on $\Delta D_{g}$, controlling for $E(\Delta D_{g}|D_{g,1})$. If $\Delta D_{g}$ is randomly assigned conditional on $D_{g,1}$, a sequential-randomization assumption as in Robins (1986), the coefficient on $\Delta D_{g}$, $\beta_{D_{1}}$, identifies a variance-weighted average of conditional averages of the treatment effect at period two (Angrist, 1998), and is therefore a “weakly causal” estimand (Blandhol et al., 2022). We propose a non-parametric two-step estimator of $\beta_{D_{1}}$.
While we consider a two-periods OLS FD regression, our results can be extended to other regressions. First, as the coefficient on $\Delta D_{g,t}$ in a “stacked” FD regression with more than two time periods and period fixed effects is a weighted average of FD coefficients estimated on pairs of consecutive time periods (Angrist, 1998), the issues we highlight also apply to stacked FD regressions. Extending the solution we propose to this case is easy: one needs to control for $E(\Delta D_{g,t}|D_{g,t-1})$ interacted with period fixed effects in the regression. Second, as the treatment coefficient in a TWFE regression is a weighted average of the treatment coefficients in first- and long-difference regressions (Ishimaru, 2022), the issues we highlight also apply to TWFE regressions. Third, as shown by de Chaisemartin and Lei (2021), the issues we highlight also apply to 2SLS FD regressions. As the reduced-form of a 2SLS is an OLS regression, our results also apply to the reduced form of 2SLS FD regressions. Therefore, the issues noted in this paper apply to a broad set of empirical applications: de Chaisemartin and D’Haultfœuille (2020) show that 19% of empirical papers published by the AER from 2010 to 2012 estimate two-way fixed effects (TWFE) or FD regressions, and many do so in designs where the treatment varies at baseline. To our knowledge, our two recommendations have not been emphasized in previous econometrics literature, and are therefore not common practice among researchers using FD or TWFE regressions. [^3] Three papers in the survey of de Chaisemartin and D’Haultfœuille (2020) estimate FD OLS regressions with one treatment variable (Algan and Cahuc, 2010; Duranton and Turner, 2011; Gentzkow et al., 2011), with a baseline treatment $D_{g,1}$ that varies, exactly as considered in this paper. None of these papers reports regressions of $\Delta D_{g}$ on $D_{g,1}$ or controls for $E(\Delta D_{g}|D_{g,1})$ or $D_{g,1}$ in the FD regression.
The issues highlighted in this paper can be avoided by assuming that there does not exist a causal model, like that in (1), where the outcome’s level is a function of the treatment’s level, and by positing directly the causal model in FD. Then, the data’s panel dimension can be ignored, and the FD regression can be analyzed as a cross-sectional regression. With a randomly-assigned treatment, a cross-sectional regression always identifies a convex combination of effects (see, e.g., Borusyak and Hull, 2024). However, the FD specification does not always come from the researcher’s belief that the causal relation between the treatment and the outcome only arises in FD. Rather, it may come from the researcher’s belief that first-differencing is necessary to net out time-invariant counfounders. Researchers ruling out a causal model in levels should make this assumption explicit, and should explain why the theory underlying their estimation delivers a causal model in FD but not in levels. [^4] Economic models underlying researchers’ estimations do not always deliver causal models in first-difference. For instance, in Section II.A of their Web Appendix, Autor et al. (2013), a paper related to that in our empirical illustration, motivate their empirical specification with a small open economy model relating “total employment in traded goods” in a US commuting zone (see page 7) to their treatment variable, rather than the evolution of total employment to the evolution of their treatment variable.
Empirical example
To illustrate our results, we apply them to the dataset of Acemoglu et al. (2016), who estimate FD regressions of the employment evolutions of US industries on the evolution of their Chinese imports penetration ratio. The paper’s Table 2 shows several FD 2SLS regressions (Columns (5) to (8)) and one stacked FD OLS regression with two first differences and three periods (Column (2)), but it does not contain any of the two-periods FD OLS regressions we consider. To ensure that our empirical application is closely aligned with our theoretical results, before considering regressions that were estimated in Acemoglu et al. (2016) we first consider an FD regression. We choose the 1999 to 2007 FD regression of US industries employment evolutions on the evolution of their Chinese imports penetration ratio, the two-periods FD regression whose coefficient is the closest to that in the stacked FD regression in the paper’s Table 2 Column (2) (-0.78 versus -0.81). Both coefficients are highly significant, suggesting a negative effect of Chinese imports on US employment.
We start by regressing $\Delta D_{g}$ on $D_{g,1}$, and we find that the two variables are strongly positively correlated: industries where imports increased the most already had a larger imports in 1999. Then, $\widehat{\beta}$ could suffer from an omitted variable bias, and here is some intuition. Assume that the US-employment response to one dollar of Chinese imports becomes more negative from 1999 to 2007 ($S_{2}<S_{1}<0$ in (1)). For instance, Chinese imports might become more substitutable to US production over time, as Chinese firms increase their value added (Chor et al., 2021). Then, an industry $h$ with high imports $D_{h,1}$ in 1999 will experience a larger employment decrease than an industry $\ell$ with low 1999 imports $D_{\ell,1}$, even if the imports of $h$ and $\ell$ do not change from 1999 to 2007. Because $S_{2}<S_{1}<0$, the employment of $h$ is more negatively affected by imports $D_{h,1}$ in 2007 than in 1999. The employment of $\ell$ is also more negatively affected by imports $D_{\ell,1}$ in 2007 than in 1999, but the change is more negative for $h$ owing to its larger baseline imports: $D_{\ell,1}<D_{h,1}\Rightarrow D_{h,1}(S_{2}-S_{1})<D_{\ell,1}(S_{2}-S_{1})$. Then, the FD regression will mistakenly attribute to $\Delta D_{g}$ decreases in US employment that would have happened without any import change, leading it to overestimate the effect of those imports. More formally, we show that even if the paths of Chinese imports $(D_{g,1},D_{g,2})$ are randomly assigned, $\widehat{\beta}$ estimates a non-convex combination of the effects of imports in 1999 and 2007, where the 1999 effect receives a weight equal to $-0.30$. Therefore, if both effects are negative but the 2007 effect is more negative, $\widehat{\beta}$ can be more negative than both effects.
Then, we turn to our alternative estimator $\widehat{\beta}_{D_{1}}$. Using a local linear kernel regression to estimate $E(\Delta D_{g}|D_{g,1})$, $\widehat{\beta}_{D_{1}}$ is 51% smaller than $\widehat{\beta}$, it is marginally significant (t-stat=-1.73), and it is significantly different from $\widehat{\beta}$ (t-stat=-2.30). Using instead a series estimator to estimate $E(\Delta D_{g}|D_{g,1})$, $\widehat{\beta}_{D_{1}}$ is 42% smaller than $\widehat{\beta}$, it is marginally significant (t-stat=-1.87), and it is marginally insignificantly different from $\widehat{\beta}$ (t-stat=-1.59).
Turning to some regressions estimated in Acemoglu et al. (2016), controlling for $E(\Delta D_{g,t}|D_{g,t-1})$ in the stacked regression shown in the paper’s Table 2 Column (2) yields a statistically insignificant coefficient, more than three times smaller than without that control. Similarly, when we consider the reduced-form of the 1999 to 2007 FD 2SLS regression in the paper’s Table 2 Column (7), controlling for the baseline instrument leads to a statistically insignificant coefficient, much smaller than without that control. Across the specifications we consider, controlling for the baseline treatment or instrument does not lead to large precision losses: $\widehat{\beta}_{D_{1}}$’s standard error is 8% smaller to 19% larger than $\widehat{\beta}$’s.
Related literature
Our Theorem 2 is related to Theorem 3 in de Chaisemartin and Lei (2021). That paper considers FD 2SLS Bartik regressions, that are sometimes justified by assuming that the FD of the shocks used to construct the Bartik instrument are as good as random, see Borusyak et al. (2022) and Section 3 of Borusyak and Hull (2023). As noted in de Chaisemartin and Lei (2021), issues with design-based justifications of FD regressions are not specific to FD Bartik regressions, which is why these ideas are now presented in this standalone paper. The current paper supersedes and extends de Chaisemartin and Lei (2021) concerning issues with design-based justifications of FD regressions. We are not working on de Chaisemartin and Lei (2021) anymore, but if we go back to it, future versions will not contain material overlapping with the current paper. More generally, our paper is also related to the literature considering design-based justifications of FD and TWFE regressions (see e.g. Athey and Imbens, 2022; Arkhangelsky et al., 2021). The connections with that literature are explained below.
2 Set-up
Two-periods panel.
One is interested in estimating the effect of a treatment on an outcome, using a panel data set, with $G$ independent and identically distributed units indexed by $g$, and with two time periods. For $t\in\{1,2\}$, let $Y_{g,t}$ and $D_{g,t}$ respectively denote the outcome and treatment of $g$ at $t$.
Assumption 1
$(Y_{g,1},Y_{g,2},D_{g,1},D_{g,2})_{g=1,...,G}$ are i.i.d.
As units are assumed i.i.d., we sometimes drop the $g$ subscript below.
First-difference regression.
Let $\Delta$ denote the FD operator. To estimate the treatment’s effect, researchers often regress $\Delta Y_{g}$ on an intercept and $\Delta D_{g}$, leading to
$\widehat{\beta}:=\frac{\sum_{g}(\Delta D_{g}-\Delta D_{.})\Delta Y_{g}}{\sum_{g}(\Delta D_{g}-\Delta D_{.})^{2}},$
where $\Delta D_{.}=1/G\sum_{g}\Delta D_{g}.$ Assuming that $(\Delta Y,\Delta D)$ has a finite second moment, under Assumption 1 the probability limit of $\widehat{\beta}$ when $G\rightarrow+\infty$ is equal to
$\beta:=\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}.$
3 The FD regression may suffer from an omitted variable bias if the treatment change and the baseline treatment are correlated.
A linear causal model in levels.
Let $Y_{g,t}(d)$ denote $g$’s potential outcome if $D_{g,t}=d$. $Y_{g,t}(0)$ is $g$’s potential outcome at $t$ if $D_{g,t}=0$, namely without treatment. Implicitly, this potential outcome notation rules out anticipation effects where $D_{g,2}$ could affect the period-one outcome, and dynamic effects where $D_{g,1}$ could affect the period-two outcome. That FD regressions rule out anticipation and dynamic effects is an important issue, different from that discussed in this paper. To simplify, we assume that the causal model is linear.
Assumption 2
For $t\in\{1,2\}$,
$Y_{t}(d)=Y_{t}(0)+S_{t}d.$ (3)
Unlike in Equation (1) in the introduction, here the slope $S_{t}$ is a random variable, that may vary across units. The slope is also indexed by $t$, as the treatment’s effect might be time-varying. Under (3), one has
$\Delta Y=Y_{2}(0)+S_{2}D_{2}-Y_{1}(0)-S_{1}D_{1}=Y_{2}(0)-Y_{1}(0)+S_{2}\times\Delta D+\Delta S\times D_{1},$ (4)
a decomposition we use repeatedly in what follows.
As good as randomly assigned first-differenced treatment.
Assumption 3
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2}).$ (5)
Assumption 3 requires that the first-differenced treatment be independent of units’ outcome evolutions without treatment, and of their period-one and period-two treatment effects.
Omitted variable bias (OVB) formula.
Theorem 1
If Assumption 3 holds,
$\beta=E(S_{2})+\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}.$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\times\Delta D)+cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}$
$=$ $E(S_{2})+\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}.$
The first equality follows from (4). The second equality follows from Assumption 3. QED.
Theorem 1 is akin to a standard omitted variable bias (OVB) formula, where $\beta$ identifies the average treatment effect at period two $E(S_{2})$, plus the OVB term $\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}$. If one is ready to assume that the treatment effect does not change over time ($S_{2}=S_{1}$), then the OVB is equal to zero and $\beta=E(S_{2})$. Similarly, if instead of Assumption 3, one is ready to make the following stronger assumption:
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2},D_{1}),$ (6)
then the OVB is also equal to zero. However, (6) implies that
$\Delta D\perp\!\!\!\perp D_{1}.$ (7)
This is a strong condition. For instance, it cannot hold if $D_{1}$ and $D_{2}$ have the same bounded support and the distribution of $\Delta D\perp\!\!\!\perp D_{1}=d$ is non-degenerate for all $d$ [^5] Letting $\underline{d}$ and $\overline{d}$ denote the support’s boundaries, the non-degenerate distribution of $\Delta D|D_{1}=\underline{d}$ cannot be the same as the non-degenerate distribution of $\Delta D|D_{1}=\overline{d}$. This is also a testable condition, that one can for instance test by regressing $\Delta D$ on $D_{1}$. Theorem 1 motivates our first recommendation: applied researchers running a first-difference regression where the baseline treatment varies should start by testing if $\Delta D$ and $D_{1}$ are correlated. If they are, and if the treatment effect varies over time ($S_{1}\neq S_{2}$), then the regression may be subject to an OVB. Intuitively, due to the correlation between $\Delta D$ and $D_{1}$, $\beta$ attributes to $\Delta D$ changes in $Y$ that would have happened without any treatment change, caused by changes in the treatment’s effect, the $\Delta S\times D_{1}$ term in (4).
Even if the treatment paths $(D_{1},D_{2})$ are randomly assigned, $\beta$ may identify a non-convex combination of $E(S_{1})$ and $E(S_{2})$.
As the OVB in the previous display is a function of $\Delta S$, one may wonder if under reasonable assumptions this OVB can simplify so that $\beta$ identifies a relatively interpretable causal effect, like a convex combination of the effects at period one and two. The following result shows this is not necessarily the case, even under the following, additional assumption:
$(D_{1},D_{2})\perp\!\!\!\perp\Delta S.$ (8)
(8) requires that the treatment paths $(D_{1},D_{2})$ be independent of units’ treatment-effect evolution $\Delta S$. Thus, (8) is more in the spirit of assuming random treatment paths than random treatment changes.
Theorem 2
If Assumptions 2-3 and (8) hold,
$\beta=\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\times\Delta D)+cov(\Delta D,\Delta S\times D_{1})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}$
$=$ $\frac{V(\Delta D)E(S_{2})+cov(\Delta D,D_{1})E(\Delta S)}{\sum_{t^{\prime}=1}^{2}V(D_{t^{\prime}})-cov(D_{1},D_{2})}$
$=$ $\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
The first equality follows from (4). The second follows from Assumption 3 and (8). QED.
Cases where the weights in Theorem 2 are guaranteed to be positive.
Theorem 2 shows that under Assumptions 2-3 and (8), $\beta$ estimates a weighted sum of the average slopes $E(S_{1})$ and $E(S_{2})$. Weights are guaranteed to be positive if the treatment is binary. [^6] To see this, note that $V(D_{1})-cov(D_{1},D_{2})=E((D_{1}-D_{2})(D_{1}-E(D_{1})))$, and $(D_{1}-D_{2})(D_{1}-E(D_{1}))\geq 0$ almost surely for binary variables. This is consistent with the results of Athey and Imbens (2022) and Arkhangelsky et al. (2021), who have shown that with a binary randomized treatment, OLS TWFE regressions always estimate a convex combination of effects. By Cauchy-Schwarz’s inequality, weights are also guaranteed to be positive if $V(D_{1})=V(D_{2})$. Weights are also guaranteed to be positive if the treatments are not serially correlated ($cov(D_{1},D_{2})=0$) or negatively correlated ($cov(D_{1},D_{2})<0$). Finally, weights are also guaranteed to be positive if the average treatment effect does not change over time ($E(S_{1})=E(S_{2})$), which is weaker than assuming that the treatment effect does not change for any unit ($S_{1}=S_{2}$). However, those positive results only hold under (8), which requires that the treatment paths, rather than their first differences, be randomly assigned. Accordingly, invoking Theorem 2 to justify a first-difference regression requires advocating for random assignment of the paths.
Cases where one of the two weights in Theorem 2 could be negative.
One of the two weights in Theorem 2 could be negative if $D_{1}$ and $D_{2}$ are non-binary, positively correlated, and have different variances. The weights in Theorem 2 can be estimated, to assess if in a given application, $\beta$ is weakly causal or not under Assumptions 2-3 and (8). Instead of Assumption 3 and (8), one could assume that
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{1}(0),Y_{2}(0),S_{1},S_{2}).$ (9)
(9) is sufficient for Assumption 3 and (8) to hold. Under (9), estimating an FD regression is not necessary anymore: to estimate the treatment’s effect, one can merely regress $Y_{1}$ on an intercept and $D_{1}$ to estimate $E(S_{1})$, and one can regress $Y_{2}$ on an intercept and $D_{2}$ to estimate $E(S_{2})$. Yet, replacing Assumption 3 and (8) by (9) does not lead to a different decomposition from that in Theorem 2: even with a randomly assigned treatment path, and under assumptions under which first-differencing is not necessary anymore, a FD regression may not estimate a convex combination of effects.
4 To avoid having an omitted variable bias, one may control for $E(\Delta D|D_{1})$ in the regression.
4.1 Identification
Let $\beta_{D_{1}}$ be the coefficient on $\Delta D$ in a regression of $\Delta Y$ on an intercept, $\Delta D$, and $E(\Delta D|D_{1})$. Let $\Delta D_{r}=\Delta D-E(\Delta D|D_{1})$ denote the residual from the CEF decomposition of $\Delta D$ with respect to $D_{1}$. We consider the following assumption.
Assumption 4
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})|D_{1}.$ (10)
Assumption 4 requires random assignment of $\Delta D$ conditional on $D_{1}$. Assumption 4 is equivalent to
$D_{2}\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})|D_{1},$
a sequential-randomization assumption as in Robins (1986).
Theorem 3
If Assumptions 2 and 4 hold,
$\beta_{D_{1}}=E\left(\frac{V(\Delta D|D_{1})}{E\left(V(\Delta D|D_{1})\right)}E(S_{2}|D_{1})\right).$
Proof
Under Assumption 4,
$cov(\Delta D_{r},Y_{2}(0)-Y_{1}(0))=cov(\Delta D_{r},S_{2})=cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))$
$=$ $cov\left(\Delta D_{r},\Delta S\times D_{1}\right)=0.$ (11)
For instance,
$cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))=$ $E(\Delta D_{r}S_{2}E(\Delta D|D_{1}))$
$=$ $E(E(\Delta D_{r}|S_{2},D_{1})S_{2}E(\Delta D|D_{1}))$
$=$ $E((E(\Delta D|S_{2},D_{1})-E(\Delta D|D_{1}))S_{2}E(\Delta D|D_{1}))$
$=$ $0$
Then,
$\beta_{D_{1}}=$ $\frac{cov(\Delta D_{r},\Delta Y)}{V(\Delta D_{r})}$
$=$ $\frac{cov(\Delta D_{r},Y_{2}(0)-Y_{1}(0))+cov(\Delta D_{r},S_{2}\times\Delta D)+cov(\Delta D_{r},\Delta S\times D_{1})}{V(\Delta D_{r})}$
$=$ $\frac{cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))+cov(\Delta D_{r},S_{2}\times\Delta D_{r})}{V(\Delta D_{r})}$
$=$ $\frac{E\left((\Delta D_{r})^{2}S_{2}\right)}{E\left((\Delta D_{r})^{2}\right)}$
$=$ $E\left(\frac{V(\Delta D|D_{1})}{E\left(V(\Delta D|D_{1})\right)}E(S_{2}|D_{1})\right).$
The first equality follows from the Frich-Waugh-Lovell theorem. The second follows from (4). The third and fourth follow from (4.1). The last follows from the law of iterated expectations and Assumption 4.
QED.
Theorem 3 shows that under Assumption 4, $\beta_{D_{1}}$ identifies a weighted average of conditional average slopes $E(S_{2}|D_{1})$. This is essentially a restatement of a well-known result in Angrist (1998) in the setting we consider. This positive result still relies on Assumption 2, which requires that the treatment effect be linear. Theorem 5 in the Appendix shows that without the linearity assumption, if the distribution of $D_{2}$ conditional on $D_{1}$ is continuous everywhere, $\beta_{D_{1}}$ identifies a weighted average, across $x$ and $d_{1}$, of $E\left(\frac{\partial Y_{2}(x)}{\partial x}\middle|D_{1}=d_{1}\right)$, the conditional average of the derivative of $x\mapsto Y_{2}(x)$ evaluated at $x$ and among units with $D_{1}=d_{1}$. If the distribution of $D_{2}$ conditional on $D_{1}$ is discrete, one can show that $\beta_{D_{1}}$ identifies a weighted average of slopes. Thus, $\beta_{D_{1}}$ remains weakly causal even if the treatment effect is non linear.
4.2 Estimation
When $D_{1}$ is a discrete variable, estimation of $E(\Delta D|D_{1})$ and then of $\beta_{D_{1}}$ is straightforward. Accordingly, in this section we focus on the case where $D_{1}$ is continuous.
Estimator.
To estimate $\beta_{D_{1}}$, one starts by estimating $E(\Delta D|D_{1})$. Then, one lets $\widehat{\beta}_{D_{1}}$ be the coefficient on $\Delta D$ in a regression of $\Delta Y$ on $\Delta D$ and $\widehat{E}(\Delta D|D_{1})$. $E(\Delta D|D_{1})$ may be estimated non-parametrically, say using local-linear kernel regressions. Alternatively, $E(\Delta D|D_{1})$ may be estimated parametrically, assuming a parametric functional-form for $E(\Delta D|D_{1})$, say a polynomial of a certain order. Then researchers can assess results’ sensitivity to the chosen functional form, by varying the polynomial order.
Influence function.
The derivation of the influence function of two-step estimators with a non-parametric first-step estimator is well understood, so we just give a heuristic derivation of the asymptotic variance of $\widehat{\beta}_{D_{1}}$, following Newey (1994). Let $\mu_{0}(D_{1})=E(\Delta D|D_{1})$. Let $\gamma_{D_{1}}=E((\Delta D-\mu_{0}(D_{1}))\Delta Y)$ and $\theta_{D_{1}}=E((\Delta D-\mu_{0}(D_{1}))^{2})$ respectively denote the numerator and denominator of $\beta_{D_{1}}$, and let $\widehat{\gamma}_{D_{1}}$ and $\widehat{\theta}_{D_{1}}$ denote the corresponding estimators. Let $Z=(D_{1},D_{2},Y_{1},Y_{2})$. Let $g(z,\mu_{0},\gamma_{D_{1}})=(\Delta d-\mu_{0})\Delta y-\gamma_{D_{1}}$ denote the score function attached to $\gamma_{D_{1}}$. The oracle influence function of $\widehat{\gamma}_{D_{1}}$ is
$\psi^{*}_{1}=(\Delta D-\mu_{0}(D_{1}))\Delta Y-\gamma_{D_{1}}.$
To find the adjustment factor to $\psi^{*}_{1}$ due to the first step estimation of $\mu_{0}(D_{1})$,
$\frac{\partial}{\partial\eta}E\left[g(Z,\mu_{0}(\eta),\gamma_{D_{1}})\right]$ $=E\left[-\dot{\mu}_{0}(D_{1},\eta)\Delta Y\right]$
$=\frac{\partial}{\partial\eta}E\left[-\mu_{0}(D_{1},\eta)\Delta Y\right],$
and thus we see that Condition (4.1) in Newey (1994) holds. Next, letting $\lambda_{1}(D_{1})=E(-\Delta Y|D_{1})$, it readily follows from the law of iterated expectations that
$E\left[-\mu_{0}(D_{1},\eta)\Delta Y\right]=$ $E\left[\mu_{0}(D_{1},\eta)\lambda_{1}(D_{1})\right],$
so Condition (4.4) in Newey (1994) holds. Thus, the adjustment factor to $\psi^{*}_{1}$ is $E(-\Delta Y|D_{1})(\Delta D-\mu_{0}(D_{1}))$. Finally, the influence function of $\widehat{\gamma}_{D_{1}}$ is
$\psi_{1}=(\Delta D-E(\Delta D|D_{1}))(\Delta Y-E(\Delta Y|D_{1}))-\gamma_{D_{1}}.$
Similarly, one can show that the influence function of $\widehat{\theta}_{D_{1}}$ is
$\psi_{2}=(\Delta D-E(\Delta D|D_{1}))^{2}-\theta_{D_{1}}.$
Finally, following standard results to linearize ratios, the influence function of $\widehat{\beta}_{D_{1}}$ is
$\psi:=\frac{\psi_{1}-\psi_{2}\beta_{D_{1}}}{\theta_{D_{1}}}.$
Letting $\widehat{\psi}$ denote a variable where all the population quantities in $\psi$ are replaced by their sample equivalents, we use the sample variance of $\widehat{\psi}$ divided by the sample size $G$ to estimate the variance of $\widehat{\beta}_{D_{1}}$.
Doubly-robust estimator?
An issue with classical two-step estimators is that tuning parameters (the bandwidths in the local-linear kernel regressions, or the polynomial order in series regressions) are not chosen in a data-driven manner. On that front, estimators solving moment conditions that are doubly-robust in the non-parametric nuisance functions offer an appealing alternative. However, in the second section of the Appendix we show that doubly-robust estimation of $\beta_{D_{1}}$ following Robinson (1988) requires that there exist stayers, namely units such that $\Delta D=0$, a condition that does not always hold in empirical applications. Moreover, this approach also requires assuming that the period-two treatment effect be mean independent of $D_{1}$, $E(S_{2}|D_{1})=E(S_{2})$, a strong assumption. An alternative would be to consider the generalization of the Augmented Inverse Propensity Weighted (AIPW) estimand of Robins et al. (1994) to continuous treatments proposed by Colangelo and Lee (2020). However, that paper proposes an estimator of the dose-response function, which converges at a slow, non-parametric rate, but does not propose a method to aggregate the dose-response function into a scalar that can be estimated at the parametric rate.
4.3 Comparison with heterogeneity-robust difference-in-differences estimands.
An alternative to $\beta_{D_{1}}$ would be to use one of the heterogeneity-robust difference-in-differences (DID) estimands that have been proposed for designs where the treatment varies at baseline (de Chaisemartin and D’Haultfœuille, 2020; de Chaisemartin and D’Haultfœuille, 2021; de Chaisemartin et al., 2022). A commonality between those DID estimands and $\beta_{D_{1}}$ is that they also control for $D_{1}$, by comparing the outcome evolutions of switchers and stayers with the same $D_{1}$. A disadvantage of the DID estimands is that they can only be used if there are stayers, a condition that does not always hold in empirical applications. An advantage of the DID estimands is that they only rely on a parallel trends assumption, and unlike $\beta_{D_{1}}$ they do not require that $D_{2}$ be independent of the treatment effect conditional on $D_{1}$.
4.4 Placebo tests.
Placebo tests are an important step in establishing the credibility of an identifying assumption in observational studies (Imbens et al., 2001; Imbens and Xu, 2024). However, placebo testing Assumption 4 is not straightforward. Assume one observes another period of data, period 0, and let $\Delta Y_{-1}$ and $\Delta D_{-1}$ respectively denote the outcome and treatment evolutions from period 00 to $1$. The coefficient on $\Delta D$ in the regression of $\Delta Y_{-1}$ on an intercept, $\Delta D$, and $E(\Delta D|D_{1})$ could differ from zero, even if Assumption 4 holds with all the unobservables switched back by one period, if $\Delta D$ is correlated to $\Delta D_{-1}$. Instead, we now propose a partly testable assumption under which a weighted average of conditional slopes is identified. Let $\beta_{D_{0},D_{1}}$ denote the coefficient on $\Delta D$ in a regression of $\Delta Y$ on an intercept, $\Delta D$, and $E(\Delta D|D_{0},D_{1})$. Let $\Delta D_{r,2}=\Delta D-E(\Delta D|D_{0},D_{1})$ denote the residual from the CEF decomposition of $\Delta D$ with respect to $(D_{0},D_{1})$. Let $\beta^{pl}_{D_{0},D_{1}}$ denote the coefficient on $\Delta D$ in a regression of $\Delta Y_{-1}$ on an intercept, $\Delta D$, and $E(\Delta D|D_{0},D_{1})$.
Assumption 5
$\Delta D\perp\!\!\!\perp(Y_{1}(0)-Y_{0}(0),Y_{2}(0)-Y_{1}(0),S_{0},S_{1},S_{2})|D_{0},D_{1}$ (12)
Assumption 5 generalizes Assumption 4 to settings with three dates, by assuming that $\Delta D$ is independent of potential outcome evolutions without treatment and treatment effects at all dates, conditional on the period zero and one treatments.
Theorem 4
If Assumptions 2 and 5 hold,
$\beta_{D_{0},D_{1}}=E\left(\frac{V(\Delta D|D_{0},D_{1})}{E\left(V(\Delta D|D_{0},D_{1})\right)}E(S_{2}|D_{0},D_{1})\right),$
and $\beta^{pl}_{D_{0},D_{1}}=0.$
Proof
The proof of the first equality is similar to the proof of the first equality in Theorem 3, and is therefore omitted. Turning to the second equality,
$\beta^{pl}_{D_{0},D_{1}}=$ $\frac{cov(\Delta D_{r,2},\Delta Y_{-1})}{V(\Delta D_{r,2})}$
$=$ $\frac{cov(\Delta D_{r,2},Y_{1}(0)-Y_{0}(0))+cov(\Delta D_{r,2},S_{1}\Delta D_{-1})+cov(\Delta D_{r,2},\Delta S_{-1}D_{0})}{V(\Delta D_{r,2})}$
$=$ $\frac{cov(\Delta D_{r,2},S_{1}D_{1})-cov(\Delta D_{r,2},S_{1}D_{0})}{V(\Delta D_{r,2})}$
$=$ $0$
QED.
5 Application
Reanalysing a two-periods OLS FD regression.
We use the panel data set constructed by Acemoglu et al. (2016), who estimate FD regressions of US industries’ employment evolutions on the evolution of their Chinese imports penetration ratio. The paper’s Table 2 shows several FD 2SLS regressions (Columns (5) to (8)) and one stacked FD OLS regression with two first differences and three periods (Column (2)), but it does not contain any of the two-periods FD OLS regressions we consider. To ensure our empirical application is closely aligned with our theoretical results, we start by revisiting one two-periods FD OLS regression, before considering regressions that were estimated in the paper. We consider the 1999 to 2007 FD regression of US industries employment evolutions on the evolution of their Chinese imports penetration ratio, the two-periods FD regression whose Chinese-imports-penetration coefficient is the closest to that in the stacked FD regression in the paper’s Table 2 Column (2). Like in the paper, all linear regressions below are weighted by industries’ 1991 employment, and standard errors are clustered at the three-digit industry level. We find that $\widehat{\beta}=-0.78$ (s.e.=$0.22$), which is very close to the coefficient in the paper’s Table 2 Column (2) ($-0.81$, s.e.=$0.16$). According to this regression, an increase in the Chinese-import-penetration ratio decreases US employment.
Testing if $\Delta D$ and $D_{1}$ are correlated.
We start by testing (7), by regressing $\Delta D$, the change in industries’ import-penetration ratio, on $D_{1}$, industries’ import-penetration ratio in 1999. The coefficient on $D_{1}$ is equal to $0.74$, and it is highly significant (s.e.=$0.16$). Thus, the industries that experienced the largest growth of their imports-penetration ratio from 1999 to 2007 already had a larger imports-penetration ratio in 1999.
Estimating the weights attached to $\beta$ in Theorem 2.
Next, we estimate the weights attached to $\beta$ in Theorem 2. We find that $\beta$ estimates a weighted sum of the average effects of import penetration in 1999 and 2007, where the 2007 average effect receives a weight equal to 1.3, while the 1999 average effect receives a weight equal to -0.3. Thus, under Assumptions 2, 3, and (8), the FD regression does not estimate a convex combination of effects, if the effect of import penetration on US employment changes from 1999 to 2007.
$E(\Delta D|D_{1})$ is non-linear.
The linearity test of Stute (1997) and Stute et al. (1998), implemented using the stute_test Stata package (de Chaisemartin et al., 2024), is very strongly rejected, with a p-value essentially equal to zero. It is only when one tests that $E(\Delta D|D_{1})$ is a polynomial of order 4 in $D_{1}$ that the test is not rejected (p-value=0.21). The stute_test package currently does not allow for weighting and to cluster standard errors, but linear regressions of $\Delta D$ on a polynomial in $D_{1}$, with the same weighting and clustering as that used by the authors, yield very similar conclusions: in a regression of $\Delta D$ on a polynomial of order 4 in $D_{1}$, all polynomial terms have t-stats larger than 5. Therefore, estimating $\beta_{D_{1}}$ cannot simply be achieved by regressing $\Delta Y$ on $\Delta D$ and $D_{1}$.
Estimation of $\beta_{D_{1}}$.
There are only five “stayers” in this application, namely industries whose Chinese-import-penetration ratio does not change over time. Therefore, a doubly-robust estimator, at least one in the spirit of Robinson (1988), cannot be used. Then, we start by estimating $E(\Delta D|D_{1})$ non-parametrically using local linear kernel regressions, with Stata’s npregress kernel command, and the command’s default choices of kernel (Epanechnikov) and bandwidth (minimizer of the integrated mean-squared error). Two industries are dropped in that estimation. Then, we regress $\Delta Y$ on $\Delta D$, controlling for $\widehat{E}(\Delta D|D_{1})$. $\widehat{\beta}_{D_{1}}$ is equal to $-0.38$, it is marginally significant (bootstrap s.e.=0.22, t-stat=-1.73), and it is significantly different from $\widehat{\beta}$ (t-stat=-2.30). [^7] When comparing $\widehat{\beta}$ and $\widehat{\beta}_{D_{1}}$, the standard error of the difference between the coefficients is computed using a bootstrap clustered at the level of three-digit industries. Using instead a series estimator to estimate $E(\Delta D|D_{1})$, with Stata’s npregress series command, leads to a point estimate of $-0.45$, which is marginally significant (s.e.=0.24, t-stat=-1.87), and marginally insignificantly different from $\widehat{\beta}$ (t-stat=-1.59). As a further robustness check, we also use a parametric polynomial model of order 4 to estimate $E(\Delta D|D_{1})$, where the choice of the polynomial order is based on the Stute tests. The coefficient on $\Delta D$ is now equal to $-0.32$ and it is marginally insignificant (s.e.=0.21, t-stat=-1.52).
Placebo test, and placebo-testable estimator.
To conduct a placebo test and obtain an estimator under a placebo-testable assumption, we estimate $E(\Delta D|D_{0},D_{1})$ using a local linear kernel regression, and we regress $\Delta Y_{-1}$ on $\Delta D$, controlling for $\widehat{E}(\Delta D|D_{0},D_{1})$. The coefficient is equal to $-0.10$ and it is insignificant (s.e. computed by a bootstrap clustered at the three-digit industry level=0.18), thus lending some credibility to Assumption 5. We also regress $\Delta Y$ on $\Delta D$, controlling for $\widehat{E}(\Delta D|D_{0},D_{1})$, thus yielding a treatment-effect estimator relying on the placebo-testable Assumption 5. The coefficient is equal to $-0.34$, and it is insignificant (bootstrap s.e.=0.28).
Re-analysis of the stacked regression in the paper’s Table 2 Column (2).
The paper’s Table 2 Column (2) contains a stacked FD regression with three periods and two first-differences, of $\Delta Y_{t}$ on $\Delta D_{t}$ and a fixed effect (FE) for the second FD. The coefficient on $\Delta D_{t}$ is equal to $-0.81$ (s.e.=$0.16$). Instead, we regress $\Delta Y_{t}$ on $\Delta D_{t}$, an FE for the second FD, $\widehat{E}(\Delta D_{t}|D_{t-1})$, and $\widehat{E}(\Delta D_{t}|D_{t-1})$ interacted with the FE for the second FD. $\widehat{E}(\Delta D_{t}|D_{t-1})$ is estimated using local linear kernel regressions. The coefficient on $\Delta D_{t}$ is small and insignificant ($-0.23$, s.e.=0.19), and significantly different from the paper’s coefficient at the 10% level (t-stat=-1.90).
Re-analysis of the reduced-form of the 2SLS regression in the paper’s Table 2 Column (7).
On top of FD OLS regressions, Acemoglu et al. (2016) also estimate FD 2SLS regressions, where $\Delta D$ is instrumented by $\Delta Z$, the change in industries’ Chinese-import-penetration ratio in eight other high-income countries. We reanalyze the reduced-form regression attached to the two-periods FD 2SLS regression in Table 2 Column (7) of the paper, namely a regression of $\Delta Y$ on $\Delta Z$ taking first differences from 1999 to 2007 as before. The coefficient on $\Delta Z$ is negative and significant ($-1.40$, s.e.=0.38). When we regress $\Delta Y$ on $\Delta Z$, controlling for $\widehat{E}(\Delta Z|Z_{1})$, estimated using a local linear kernel regression, the coefficient is smaller than before and insignificant ($-0.36$, bootstrap s.e.=0.35). The difference with the paper’s reduced-form coefficient is insignificant (t-stat=-1.43).
Table 1: First-difference regressions of employment on Chinese imports
Time periods Indep var Controls? Estimated in paper? Coefficient s.e. N
1999 to 2007 $\Delta D$ No No $-0.78$ $0.22$ 392
1999 to 2007 $\Delta D$ $E(\Delta D|D_{1})$, kernel No $-0.38$ $0.22$ 390
1999 to 2007 $\Delta D$ $E(\Delta D|D_{1})$, series No $-0.45$ $0.24$ 392
1999 to 2007 $\Delta D$ $D_{1}$, $D_{1}^{2}$, $D_{1}^{3}$, $D_{1}^{4}$ No $-0.32$ $0.21$ 392
Stacked $\Delta D_{t}$ No Table 2 Column (2) $-0.81$ $0.16$ 784
Stacked $\Delta D_{t}$ $E(\Delta D_{t}|D_{t-1})$, kernel No $-0.23$ $0.19$ 782
1999 to 2007 $\Delta Z$ No RF of Table 2 Column (7) $-1.40$ $0.38$ 392
1999 to 2007 $\Delta Z$ $E(\Delta Z|Z_{1})$, kernel No $-0.36$ $0.35$ 392
The table shows regressions of US industries employment evolutions $\Delta Y$ on the evolution of their Chinese import-penetration ratio $\Delta D$, using the data of Acemoglu et al. (2016). Regressions are weighted by industries’ 1991 employment, and standard errors are clustered on 135 three-digit industries, as in the paper. The first line shows a FD regression from 1999 to 2007. The second, third, and fourth lines are similar, but control for $E(\Delta D|D_{1})$, estimated using a local-linear kernel regression (second line), a series estimator (third line), and a fourth-order polynomial (fourth line). The fifth line re-estimates the stacked regression in the paper’s Table 2 Column (2). The sixth line is similar to the fifth, but controls for $E(\Delta D_{t}|D_{t-1})$, estimated using local-linear kernel regressions. The seventh line re-estimates the reduced-form of the paper’s 2SLS regression in Table 2 Column (7). The eighth line is similar, but controls for $E(\Delta Z|Z_{1})$, estimated using a local-linear kernel regression.
6 Conclusion
Under a causal model in levels, first-difference regressions may suffer from an omitted variable bias if the treatment change $\Delta D$ is correlated to the baseline treatment $D_{1}$, and the treatment effect changes over time. A simple fix is to control for $E(\Delta D|D_{1})$.