### v7

Abstract
We consider treatment-effect estimation using a first-difference regression of an outcome evolution $\Delta Y$ on a treatment evolution $\Delta D$. Under a causal model in levels with a time-varying effect, the regression residual is a function of the period-one treatment $D_{1}$. Then, researchers should test if $\Delta D$ and $D_{1}$ are correlated: if they are, the regression may suffer from an omitted variable bias. To solve it, researchers may control nonparametrically for $E(\Delta D|D_{1})$. We use our results to revisit first-difference regressions estimated on the data of Acemoglu et al. (2016), who study the effect of imports from China on US employment. $\Delta D$ and $D_{1}$ are strongly correlated, thus implying that first-difference regressions may be biased if the effect of Chinese imports changes over time. The coefficient on $\Delta D$ is no longer significant when controlling for $E(\Delta D|D_{1})$.
1 Introduction
Assume that one seeks to estimate a treatment’s effect with a two-periods panel, where units’ exposure to treatment varies at period one. If one worries that time-invariant confounders could bias a naive regression of the outcome on the treatment, one may use a first-difference (FD) regression of the outcome evolution $\Delta Y_{g}$ on the treatment evolution $\Delta D_{g}$. Let $\widehat{\beta}$ denote the coefficient on $\Delta D_{g}$. To justify $\widehat{\beta}$, two identifying assumptions are available in the econometrics literature. First, one may invoke a model-based parallel-trends assumption, but that assumption may not be sufficient to ensure that $\widehat{\beta}$ estimates a convex combination of unit-and-period specific effects (de Chaisemartin and D’Haultfœuille, 2020; de Chaisemartin and Lei, 2021). Second, one may assume that $\Delta D_{g}$ is as good as randomly assigned, namely uncorrelated to the effect of $\Delta D_{g}$, and to the residual of a regression of $\Delta Y_{g}$ on $\Delta D_{g}$. We consider this second, design-based justification.
Assume that $Y_{g,t}$, the outcome of unit $g$ at $t$, is generated by a two-way fixed effects model with a linear effect of the treatment $D_{g,t}$:
$Y_{g,t}=\theta_{g}+\alpha_{t}+S_{t}D_{g,t}+u_{g,t}.$ (1)
$S_{t}$ is the effect of increasing the treatment by one unit on the outcome at period $t$. To simplify, we assume for now that the treatment effect is constant across units, but we allow it to vary over time. First-differencing (1), and letting $v_{g}=D_{g,1}(S_{2}-S_{1})+\Delta u_{g}$,
$\Delta Y_{g}=$ $\theta_{g}+\alpha_{2}+S_{2}D_{g,2}+u_{g,2}-(\theta_{g}+\alpha_{1}+S_{1}D_{g,1}+u_{g,1})$
$=$ $\Delta\alpha+S_{2}\Delta D_{g}+D_{g,1}(S_{2}-S_{1})+\Delta u_{g}$
$=$ $\Delta\alpha+S_{2}\Delta D_{g}+v_{g}.$ (2)
Then, under (1), the residual of a regression of $\Delta Y_{g}$ on $\Delta D_{g}$ is $v_{g}=D_{g,1}(S_{2}-S_{1})+\Delta u_{g}$, and $\Delta D_{g}$ has to be uncorrelated to $v_{g}$ for the regression not to suffer from an omitted variable bias (OVB). But if the treatment effect changes over time ($S_{2}\neq S_{1}$), $v_{g}$ is a function of $D_{g,1}$, so having that $\Delta D_{g}$ and $v_{g}$ are uncorrelated essentially requires that $\Delta D_{g}$ and $D_{g,1}$ are independent. This a strong condition: for instance, it cannot hold if $D_{g,1}$ and $D_{g,2}$ have the same bounded support, or if $D_{g,2}$ follows a stationary AR(1) process. This is also a testable condition, and our first recommendation is that researchers invoking random assignment to justify $\widehat{\beta}$ start by regressing $\Delta D_{g}$ on $D_{g,1}$. This is akin to a standard balancing check. But while usually, which variables to include in balancing checks is open to interpretation, here the residual is a function of $D_{g,1}$, so $D_{g,1}$ needs to be included.
Here is some intuition on the OVB: due to the correlation between $\Delta D_{g}$ and $D_{g,1}$, $\widehat{\beta}$ attributes to $\Delta D_{g}$ changes in $Y_{g,t}$ that would have happened without any treatment change, that are caused by changes in the treatment’s effect, the $D_{g,1}(S_{2}-S_{1})$ term in $v_{g}$. Below, we give further intuition in the context of our empirical example. As the OVB is a function of the treatment-effect’s evolution $S_{2}-S_{1}$, one may wonder if under reasonable assumptions this OVB simplifies, so that $\widehat{\beta}$ eventually estimates a convex combination of $S_{1}$ and $S_{2}$. We show that this is not necessarily the case: even if the treatment paths $(D_{g,1},D_{g,2})$ are randomly assigned, which is stronger than assuming that $\Delta D_{g}$ is randomized, $\widehat{\beta}$ may estimate a non-convex combination of effects. With randomly-assigned paths, estimating an FD regression is not necessary: one can merely regress $Y_{g,1}$ (resp. $Y_{g,2}$) on an intercept and $D_{g,1}$ (resp. $D_{g,2}$) to estimate the treatment effect at period one (resp. two). Thus, our result shows that under assumptions under which first-differencing is unnecessary, a FD regression may not estimate a convex combination of effects: there are instances where first-differencing can do more harm than good.
If $\Delta D_{g}$ and $D_{g,1}$ are correlated and researchers are not willing to assume a constant effect over time, our second recommendation is that they may estimate the treatment’s effect by regressing $\Delta Y_{g}$ on $\Delta D_{g}$, controlling non-parametrically for $E(\Delta D_{g}|D_{g,1})$. If $D_{g,2}$ is randomly assigned conditional on $D_{g,1}$, a sequential-randomization assumption related to that in Robins (1986), the coefficient on $\Delta D_{g}$, $\beta_{D_{1}}$, identifies a variance-weighted average of conditional averages of the treatment effect at period two (Angrist, 1998), and is therefore a “weakly causal” estimand (Blandhol et al., 2022). To estimate $\beta_{D_{1}}$, we follow Chernozhukov et al. (2018) and use a Double-Debiased Machine Learning (DDML) estimator.
While we consider a two-periods OLS FD regression, our results can be extended to other regressions. First, as the coefficient on $\Delta D_{g,t}$ in a “stacked” FD regression with more than two time periods and period fixed effects is a weighted average of FD coefficients estimated on pairs of consecutive time periods (Angrist, 1998), the issues we highlight also apply to stacked FD regressions. Extending the solution we propose to this case is easy: one needs to control for $E(\Delta D_{g,t}|D_{g,t-1})$ interacted with period fixed effects in the regression. Second, as the treatment coefficient in a two-way fixed effects (TWFE) regression is a weighted average of the treatment coefficients in first- and long-difference regressions (Ishimaru, 2022), the issues we highlight also apply to TWFE regressions. Third, as shown by de Chaisemartin and Lei (2021), the issues we highlight also apply to 2SLS FD regressions. As the reduced-form of a 2SLS is an OLS regression, our results also apply to the reduced form of 2SLS FD regressions. Therefore, the issues noted in this paper apply to a broad set of empirical applications: de Chaisemartin and D’Haultfœuille (2020) show that 19% of empirical papers published by the AER from 2010 to 2012 estimate TWFE or FD regressions, and many do so in designs where the treatment varies at baseline. Parallel-trends is typically not sufficient to ensure that those regressions estimate a convex combination of effects, and we show that invoking random assignment of treatment changes to justify those regressions also comes with some caveats. To our knowledge, our two recommendations have not been emphasized in previous econometrics literature, and are therefore not common practice among researchers using FD or TWFE regressions.
The issues highlighted in this paper can be avoided by assuming there does not exist a causal model, like that in (1), where the outcome’s level is a function of the treatment’s level, and by positing directly the causal model in FD. Then, the data’s panel dimension can be ignored, and the FD regression can be analyzed as a cross-sectional regression. With a randomly-assigned treatment, a cross-sectional regression without covariates always identifies a convex combination of effects (see, e.g., Borusyak and Hull, 2024). However, the FD specification does not always come from the researcher’s belief that the causal model only arises in FD. Rather, it may come from the researcher’s belief that first-differencing is necessary to net out time-invariant counfounders. Researchers ruling out a model in levels should make this assumption explicit, and should explain why the theory underlying their estimation delivers a model in FD but not in levels. [^4] Economic models underlying researchers’ estimations do not always deliver causal models in first-difference. For instance, in Section II.A of their Web Appendix, Autor et al. (2013), a paper related to that in our empirical illustration, motivate their empirical specification with a small open economy model relating “total employment in traded goods” in a US commuting zone (see page 7) to their treatment variable, rather than the evolution of total employment to the evolution of their treatment variable.
Empirical example
To illustrate our results, we apply them to the dataset of Acemoglu et al. (2016), who estimate FD regressions of the employment evolutions of US industries on the evolution of their Chinese imports penetration ratio. The paper’s Table 2 shows several FD 2SLS regressions (Columns (5) to (8)) and one stacked FD OLS regression with two first differences and three periods (Column (2)), but it does not contain any of the two-periods FD OLS regressions we consider. To ensure that our empirical application is closely aligned with our theoretical results, before considering regressions that were estimated in the paper, we first consider a two-periods FD regression. We choose the 1999 to 2007 FD regression of US industries employment evolutions on the evolution of their Chinese imports, the two-periods FD regression whose coefficient is the closest to that in the stacked FD regression in the paper’s Table 2 Column (2) (-0.78 versus -0.81). Both coefficients are highly significant, suggesting a negative effect of Chinese imports on US employment.
First, we follow de Chaisemartin and D’Haultfœuille (2020), and decompose the FD regression under the parallel-trends assumption. Under that assumption, $\widehat{\beta}$ estimates a weighted sum of industry×$\times$×year specific effects of Chinese imports on employment, where half of the effects are weighted negatively, and negative weights sum to $-0.35$. Thus, parallel-trends is not sufficient to ensure that $\widehat{\beta}$ is weakly causal, which motivates investigating whether $\widehat{\beta}$ is weakly causal under the assumption that $\Delta D$ is as good as random.
We start our investigation by regressing $\Delta D_{g}$ on $D_{g,1}$, and we find that the two variables are strongly positively correlated: industries where imports increased the most already had larger imports in 1999. Then, $\widehat{\beta}$ could suffer from an omitted variable bias, and here is some intuition. Assume that the US-employment response to one dollar of Chinese imports becomes more negative from 1999 to 2007 ($S_{2}<S_{1}<0$ in (1)). For instance, Chinese imports might become more substitutable to US production over time, as Chinese firms increase their value added (Chor et al., 2021). Then, consider an industry $h$ with high imports $D_{h,1}$ in 1999, and an industry $\ell$ with low 1999 imports $D_{\ell,1}$. Because $S_{2}<S_{1}<0$, the employment of $h$ is more negatively affected by $D_{h,1}$ imports in 2007 than in 1999. The employment of $\ell$ is also more negatively affected by $D_{\ell,1}$ imports in 2007 than in 1999, but the change in the effect is more negative for $h$ owing to its larger baseline imports: $D_{\ell,1}<D_{h,1}\Rightarrow D_{h,1}(S_{2}-S_{1})<D_{\ell,1}(S_{2}-S_{1})$. Then, even if industries’ imports do not change from 1999 to 2007, industry $h$ will experience a larger employment decrease from 1999 to 2007 than industry $\ell$. Therefore, due to the correlation between $\Delta D_{g}$ on $D_{g,1}$, the FD regression will mistakenly attribute to $\Delta D_{g}$ decreases in employment that would have happened without any import change, leading it to overestimate the effect of those imports. More formally, we show that even if the paths of Chinese imports $(D_{g,1},D_{g,2})$ are randomly assigned, $\widehat{\beta}$ estimates a non-convex combination of the effects of imports in 1999 and 2007, where the 1999 effect receives a weight equal to $-0.30$. Then, if both effects are negative but the 2007 effect is more negative, $\widehat{\beta}$ can be more negative than both effects.
Then, we turn to our alternative estimator $\widehat{\beta}_{D_{1}}$. Controlling for $E(\Delta D_{g}|D_{g,1})$ nonparametrically using a DDML estimator yields a coefficient 25% smaller than $\widehat{\beta}$, with a standard error that is almost twice larger. Therefore, $\widehat{\beta}_{D_{1}}$ is not significantly different from zero. We also find evidence of differential pre-trends: conditional on lagged treatments, $\Delta D_{g}$ is significantly correlated with industries’ 1991-to-1999 employment evolutions. Turning to some regressions estimated in Acemoglu et al. (2016), controlling for $E(\Delta D_{g,t}|D_{g,t-1})$ in the stacked regression shown in the paper’s Table 2 Column (2) yields an insignificant coefficient, more than 33% smaller than in the paper, and with a standard error around 2.5 times larger. Similarly, when we consider the reduced-form of the 1999 to 2007 FD 2SLS regression in the paper’s Table 2 Column (7), controlling for the baseline instrument yields an insignificant coefficient, more than 40% smaller than in the paper and with a standard error around 2.5 times larger.
Related literature
Our paper is related to the literature considering design-based justifications of FD and TWFE regressions (see e.g. Athey and Imbens, 2022; Arkhangelsky et al., 2021). The connections with those papers are explained below. At this stage, we note that our Theorem 2 is related to Theorem 3 in de Chaisemartin and Lei (2021). That paper considers FD 2SLS Bartik regressions, that are sometimes justified by assuming that the FD of the shocks used to construct the Bartik instrument are as good as random, see Borusyak et al. (2022) and Section 3 of Borusyak and Hull (2023). [^5] In de Chaisemartin and Lei (2021), when we revisit Autor et al. (2013), the same empirical application as Borusyak et al. (2022), we find that the FD of the shocks is very strongly positively correlated to the baseline shocks, so the issues we highlight in this paper also apply to that empirical example. As noted in de Chaisemartin and Lei (2021), issues with design-based justifications of FD regressions are not specific to FD Bartik regressions, which is why these ideas are now presented in this standalone paper. The current paper supersedes and extends de Chaisemartin and Lei (2021) concerning issues with design-based justifications of FD regressions.
2 Set-up
Two-periods panel.
One is interested in estimating the effect of a treatment on an outcome, using a panel data set, with $G$ independent and identically distributed units indexed by $g$, and with two time periods. For $t\in\{1,2\}$, let $Y_{g,t}$ and $D_{g,t}$ respectively denote the outcome and treatment of $g$ at $t$.
Assumption 1
$(Y_{g,1},Y_{g,2},D_{g,1},D_{g,2})_{g=1,...,G}$ are i.i.d.
As units are assumed i.i.d., we sometimes drop the $g$ subscript below.
First-difference regression.
Let $\Delta$ denote the FD operator. To estimate the treatment’s effect, researchers often regress $\Delta Y_{g}$ on an intercept and $\Delta D_{g}$, leading to
$\widehat{\beta}:=\frac{\sum_{g}(\Delta D_{g}-\overline{\Delta D})\Delta Y_{g}}{\sum_{g}(\Delta D_{g}-\overline{\Delta D})^{2}},$
where $\overline{\Delta D}=1/G\sum_{g}\Delta D_{g}.$ Assuming that $(\Delta Y,\Delta D)$ has a finite second moment, under Assumption 1 the probability limit of $\widehat{\beta}$ when $G\rightarrow+\infty$ is equal to
$\beta:=\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}.$
3 The FD regression may suffer from an omitted variable bias if the treatment change and the baseline treatment are correlated.
A linear causal model in levels.
Let $Y_{g,t}(d)$ denote $g$’s potential outcome if $D_{g,t}=d$. $Y_{g,t}(0)$ is $g$’s potential outcome at $t$ if $D_{g,t}=0$, namely without treatment. Implicitly, this potential outcome notation rules out anticipation effects where $D_{g,2}$ could affect the period-one outcome, and dynamic effects where $D_{g,1}$ could affect the period-two outcome. That FD regressions rule out anticipation and dynamic effects is an important issue, different from that discussed in this paper. To simplify, we assume that the causal model is linear.
Assumption 2
For $t\in\{1,2\}$,
$Y_{t}(d)=Y_{t}(0)+S_{t}d.$ (3)
Unlike in Equation (1) in the introduction, here the slope $S_{t}$ is a random variable, that may vary across units. The slope is also indexed by $t$, as the treatment’s effect might be time-varying. Under (3), one has
$\Delta Y=Y_{2}(0)+S_{2}D_{2}-Y_{1}(0)-S_{1}D_{1}=Y_{2}(0)-Y_{1}(0)+S_{2}\times\Delta D+\Delta S\times D_{1},$ (4)
a decomposition we use repeatedly in what follows.
As good as randomly assigned first-differenced treatment.
Assumption 3
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2}).$ (5)
Assumption 3 requires that the first-differenced treatment be independent of units’ outcome evolutions without treatment, and of their period-one and period-two treatment effects. It is stronger than a parallel-trends assumption, that typically would only require that $Y_{2}(0)-Y_{1}(0)$ is independent or mean independent of $\Delta D$.
Omitted variable bias (OVB) formula.
Theorem 1
If Assumption 3 holds,
$\beta=E(S_{2})+\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}.$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\times\Delta D)+cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}$
$=$ $E(S_{2})+\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}.$
The first equality follows from (4). The second equality follows from Assumption 3. QED.
Theorem 1 is akin to a standard omitted variable bias (OVB) formula. It shows that $\beta$ identifies the average treatment effect at period two $E(S_{2})$, plus the OVB term
$\frac{cov(\Delta D,\Delta S\times D_{1})}{V(\Delta D)}.$
If one is ready to assume that the treatment effect does not change over time ($S_{2}=S_{1}$), then the OVB is equal to zero and $\beta=E(S_{2})$. Similarly, if instead of Assumption 3, one is ready to make the following stronger assumption:
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2},D_{1}),$ (6)
then the OVB is also equal to zero. However, (6) implies that
$\Delta D\perp\!\!\!\perp D_{1}.$ (7)
(7) is a strong condition. For instance, it cannot hold if $D_{1}$ and $D_{2}$ have the same bounded support and the distribution of $\Delta D|D_{1}=d$ is non-degenerate for all $d$ [^6] Letting $\underline{d}$ and $\overline{d}$ denote the support’s boundaries, the non-degenerate distribution of $\Delta D|D_{1}=\underline{d}$ cannot be the same as the non-degenerate distribution of $\Delta D|D_{1}=\overline{d}$. If $D_{2}$ follows an AR(1) process ($D_{2}=\lambda_{0}+\lambda_{1}D_{1}+\epsilon$), (7) can only hold in the knife-edge case where $\lambda_{1}=1$, meaning that the process is non-stationary. (7) is also testable, for instance by regressing $\Delta D$ on $D_{1}$. Theorem 1 motivates our first recommendation: applied researchers running a first-difference regression where the baseline treatment varies should start by testing if $\Delta D$ and $D_{1}$ are correlated. If they are, and if the treatment effect varies over time ($S_{1}\neq S_{2}$), then the regression may be subject to an OVB.
Even if the treatment paths $(D_{1},D_{2})$ are randomly assigned, $\beta$ may identify a non-convex combination of $E(S_{1})$ and $E(S_{2})$.
As the OVB in the previous display is a function of $\Delta S$, one may wonder if under reasonable assumptions this OVB can simplify so that $\beta$ identifies a relatively interpretable causal effect, like a convex combination of the effects at period one and two. The following result shows this is not necessarily the case, even under the following assumption:
$(D_{1},D_{2})\perp\!\!\!\perp(Y_{1}(0),Y_{2}(0),S_{1},S_{2}).$ (8)
(8) requires that the treatment paths $(D_{1},D_{2})$ be independent of units’ period-one and period-two untreated potential outcomes and treatment effects. Thus, (8) assumes random treatment paths, which is stronger than Assumption 3.
Theorem 2
If Assumption 2 and (8) hold,
$\beta=\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
Proof
$\frac{cov(\Delta D,\Delta Y)}{V(\Delta D)}=$ $\frac{cov(\Delta D,Y_{2}(0)-Y_{1}(0))+cov(\Delta D,S_{2}\times\Delta D)+cov(\Delta D,\Delta S\times D_{1})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}$
$=$ $\frac{V(\Delta D)E(S_{2})+cov(\Delta D,D_{1})E(\Delta S)}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}$
$=$ $\sum_{t=1}^{2}\frac{V(D_{t})-cov(D_{1},D_{2})}{\sum_{t^{\prime}=1}^{2}[V(D_{t^{\prime}})-cov(D_{1},D_{2})]}E(S_{t}).$
The first equality follows from (4). The second follows from (8). QED.
Cases where the weights in Theorem 2 are guaranteed to be positive.
Theorem 2 shows that under Assumption 2 and (8), $\beta$ estimates a weighted sum of the average slopes $E(S_{1})$ and $E(S_{2})$. Weights are guaranteed to be positive if the treatment is binary. [^7] To see this, note that $V(D_{1})-cov(D_{1},D_{2})=E((D_{1}-D_{2})(D_{1}-E(D_{1})))$, and $(D_{1}-D_{2})(D_{1}-E(D_{1}))\geq 0$ almost surely for binary variables. This is consistent with the results of Athey and Imbens (2022) and Arkhangelsky et al. (2021), who have shown that with a binary randomized treatment, OLS TWFE regressions always estimate a convex combination of effects. By Cauchy-Schwarz’s inequality, weights are also guaranteed to be positive if $V(D_{1})=V(D_{2})$. Weights are also guaranteed to be positive if the treatments are not serially correlated ($cov(D_{1},D_{2})=0$) or negatively correlated ($cov(D_{1},D_{2})<0$). Finally, weights are also guaranteed to be positive if the average treatment effect does not change over time ($E(S_{1})=E(S_{2})$), which is weaker than assuming that the treatment effect does not change for any unit ($S_{1}=S_{2}$). However, those positive results only hold under (8), which requires that the treatment paths, rather than their first differences, be randomly assigned. Accordingly, invoking Theorem 2 to justify a first-difference regression requires advocating for random assignment of the paths.
Cases where one of the two weights in Theorem 2 could be negative.
One of the two weights in Theorem 2 could be negative if $D_{1}$ and $D_{2}$ are non-binary, positively correlated, and have different variances. The weights in Theorem 2 can be estimated, to assess if in a given application, $\beta$ is weakly causal or not under Assumption 2 and (8).
First-differencing can do more harm than good.
Under (8), estimating an FD regression is not necessary: one can merely regress $Y_{1}$ on an intercept and $D_{1}$ to estimate $E(S_{1})$, and one can regress $Y_{2}$ on an intercept and $D_{2}$ to estimate $E(S_{2})$. Thus, Theorem 2 shows that under assumptions under which first-differencing is unnecessary, an FD regression may not estimate a convex combination of effects. This shows that there exists instances where first-differencing can do more harm than good.
4 To avoid having an omitted variable bias, one may control for $E(\Delta D|D_{1})$ in the regression.
4.1 Identification
Let $\beta_{D_{1}}$ be the coefficient on $\Delta D$ in a regression of $\Delta Y$ on an intercept, $\Delta D$, and $E(\Delta D|D_{1})$. Let $\Delta D_{r}=\Delta D-E(\Delta D|D_{1})$ denote the residual from the CEF decomposition of $\Delta D$ with respect to $D_{1}$. We consider the following assumption.
Assumption 4
$\Delta D\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})|D_{1}.$ (9)
Assumption 4 requires random assignment of $\Delta D$ conditional on $D_{1}$. It is equivalent to
$D_{2}\perp\!\!\!\perp(Y_{2}(0)-Y_{1}(0),S_{1},S_{2})|D_{1},$
a sequential-randomization assumption in the spirit of that in Robins (1986), except that Assumption 4 does not require that randomization is conditional on the lagged outcome.
Theorem 3
If Assumptions 2 and 4 hold,
$\beta_{D_{1}}=E\left(\frac{V(\Delta D|D_{1})}{E\left(V(\Delta D|D_{1})\right)}E(S_{2}|D_{1})\right).$
Proof
Under Assumption 4,
00 0 $=cov(\Delta D_{r},Y_{2}(0)-Y_{1}(0))=cov(\Delta D_{r},S_{2})=cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))$
$=$ $cov\left(\Delta D_{r},\Delta S\times D_{1}\right).$ (10)
For instance,
$cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))=$ $E(\Delta D_{r}S_{2}E(\Delta D|D_{1}))$
$=$ $E(E(\Delta D_{r}|S_{2},D_{1})S_{2}E(\Delta D|D_{1}))$
$=$ $E((E(\Delta D|S_{2},D_{1})-E(\Delta D|D_{1}))S_{2}E(\Delta D|D_{1}))$
$=$ $0$
Then,
$\beta_{D_{1}}=$ $\frac{cov(\Delta D_{r},\Delta Y)}{V(\Delta D_{r})}$ (11)
$=$ $\frac{cov(\Delta D_{r},Y_{2}(0)-Y_{1}(0))+cov(\Delta D_{r},S_{2}\times\Delta D)+cov(\Delta D_{r},\Delta S\times D_{1})}{V(\Delta D_{r})}$
$=$ $\frac{cov(\Delta D_{r},S_{2}E(\Delta D|D_{1}))+cov(\Delta D_{r},S_{2}\times\Delta D_{r})}{V(\Delta D_{r})}$
$=$ $\frac{E\left((\Delta D_{r})^{2}S_{2}\right)}{E\left((\Delta D_{r})^{2}\right)}$
$=$ $E\left(\frac{V(\Delta D|D_{1})}{E\left(V(\Delta D|D_{1})\right)}E(S_{2}|D_{1})\right).$
The first equality follows from the Frich-Waugh-Lovell theorem. The second follows from (4). The third and fourth follow from (4.1). The last follows from the law of iterated expectations and Assumption 4.
QED.
Theorem 3 shows that under Assumption 4, $\beta_{D_{1}}$ identifies a weighted average of conditional average slopes $E(S_{2}|D_{1})$. This is essentially a restatement of a well-known result in Angrist (1998) in the setting we consider. This positive result still relies on Assumption 2, which requires that the treatment effect be linear. Theorem 5 in the Appendix shows that without the linearity assumption, if the distribution of $D_{2}$ conditional on $D_{1}$ is continuous everywhere, $\beta_{D_{1}}$ identifies a weighted average, across $x$ and $d_{1}$, of $E\left(\frac{\partial Y_{2}(x)}{\partial x}\middle|D_{1}=d_{1}\right)$, the conditional average of the derivative of $x\mapsto Y_{2}(x)$ evaluated at $x$ and among units with $D_{1}=d_{1}$. If the distribution of $D_{2}$ conditional on $D_{1}$ is discrete, one can show that $\beta_{D_{1}}$ identifies a weighted average of slopes. Thus, $\beta_{D_{1}}$ remains weakly causal even if the treatment effect is non linear.
4.2 Estimation
When $D_{1}$ is a discrete variable, estimation of $E(\Delta D|D_{1})$ and then of $\beta_{D_{1}}$ is straightforward. Accordingly, in this section we focus on the case where $D_{1}$ is continuous.
For estimation purposes, we first note that (11) is equivalent to
$\beta_{D_{1}}=\frac{cov\left(\Delta Y_{r},\Delta D_{r}\right)}{V\left(\Delta D_{r}\right)},$ (12)
where $\Delta Y_{r}=\Delta Y-E\left(\left.\Delta Y\right|D_{1}\right)$. Equation (12) is convenient since it implies that $\beta_{D_{1}}$ is identified by a Neyman-orthogonal moment, locally robust to both $E\left(\left.\Delta D\right|D_{1}\right)$ and $E\left(\left.\Delta Y\right|D_{1}\right)$ (Chernozhukov et al., 2018). This implies that we can estimate these conditional expectations through nonparametric flexible procedures (e.g., machine learning) and inference on (12) will be valid under more general conditions than plug-in approaches (using, e.g., (11)), as discussed by Chernozhukov et al. (2022).
Estimator.
To estimate $\beta_{D_{1}}$, we shall use the Double-Debiased Machine Learning (DDML) procedure of Chernozhukov et al. (2018). Its implementation requires cross-fitting. [^8] Cross-fitting has a long tradition in the semiparametric literature; see, e.g., Bickel (1982), Klaassen (1987), van der Vaart (1998), Robins et al. (2008), and Zheng and van der Laan (2010). To this end, we need to randomly partition the sample into $L$ folds, $I_{1},\cdots,I_{L}$, of the same size. Let $I^{c}_{\ell}$ be the complement of $I_{\ell}$ in $\left\{1,2\cdots,n\right\}$. To simplify our discussion, let $\eta\left(D_{1}\right)=E(\Delta D|D_{1})$ and $\gamma\left(D_{1}\right)=E(\Delta Y|D_{1})$. We conduct estimation of $\beta_{D_{1}}$ as follows:
Step 1: Estimate $\eta$ and $\gamma$ using $I^{c}_{\ell}$, obtaining cross-fitted estimators $\hat{\eta}_{\ell}$ and $\hat{\gamma}_{\ell}$, for each $\ell=1,2,\cdots,L$.
Step 2: Construct residuals $\widehat{\Delta D}_{r,i}=\Delta D_{i}-\hat{\eta}_{\ell}(D_{1i})$ and $\widehat{\Delta Y}_{r,i}=\Delta Y_{i}-\hat{\gamma}_{\ell}(D_{1i})$, for each $i\in I_{\ell}$, $\ell=1,2,\cdots,L$.
Step 3: Run OLS of $\widehat{\Delta Y}_{r}$ on $\widehat{\Delta D}_{r}$. The corresponding slope is our estimator $\widehat{\beta}_{D_{1}}$.
The asymptotic properties of the previous estimator follows by a straightforward application of Theorems 3.1-3.2 of Chernozhukov et al. (2018).
4.3 Comparison with heterogeneity-robust difference-in-differences estimands.
An alternative to $\beta_{D_{1}}$ would be to use one of the heterogeneity-robust difference-in-differences (DID) estimands that have been proposed for designs where the treatment varies at baseline (de Chaisemartin and D’Haultfœuille, 2020; de Chaisemartin and D’Haultfœuille, 2021; de Chaisemartin et al., 2022). A commonality between those DID estimands and $\beta_{D_{1}}$ is that they also control for $D_{1}$: they compare the outcome evolutions of units whose treatment changes (switchers) and units whose treatment does not change (stayers) with the same $D_{1}$. A disadvantage of the DID estimands is that they can only be used if there are stayers, a condition that does not always hold in empirical applications. An advantage of the DID estimands is that they only rely on a parallel-trends assumption: unlike $\beta_{D_{1}}$ they do not require that $D_{2}$ be independent of the treatment effect conditional on $D_{1}$.
4.4 Placebo tests.
Placebo tests are an important step in establishing the credibility of an identifying assumption in observational studies (Imbens et al., 2001; Imbens and Xu, 2024). However, placebo testing Assumption 4 is not straightforward. Assume one observes another period of data, period 0, and let $\Delta Y_{-1}$, $\Delta D_{-1}$, and $\Delta S_{-1}$ respectively denote the outcome, treatment, and treatment-effect evolutions from period 00 to $1$. The coefficient on $\Delta D$ in the regression of $\Delta Y_{-1}$ on an intercept, $\Delta D$, and $E(\Delta D|D_{1})$ could differ from zero, even if Assumption 4 holds with all the unobservables switched back by one period, if $\Delta D$ is correlated to $\Delta D_{-1}$. Instead, we now propose a partly testable assumption under which a weighted average of conditional slopes is identified. Let $\beta_{D_{0},D_{1}}$ denote the coefficient on $\Delta D$ in a regression of $\Delta Y$ on an intercept, $\Delta D$, and $E(\Delta D|D_{0},D_{1})$. Let $\Delta D_{r,2}=\Delta D-E(\Delta D|D_{0},D_{1})$ denote the residual from the CEF decomposition of $\Delta D$ with respect to $(D_{0},D_{1})$. Let $\beta^{pl}_{D_{0},D_{1}}$ denote the coefficient on $\Delta D$ in a regression of $\Delta Y_{-1}$ on an intercept, $\Delta D$, and $E(\Delta D|D_{0},D_{1})$.
Assumption 5
$\Delta D\perp\!\!\!\perp(Y_{1}(0)-Y_{0}(0),Y_{2}(0)-Y_{1}(0),S_{0},S_{1},S_{2})|D_{0},D_{1}$ (13)
Assumption 5 generalizes Assumption 4 to settings with three dates, by assuming that $\Delta D$ is independent of potential outcome evolutions without treatment and treatment effects at all dates, conditional on the period zero and one treatments.
Theorem 4
If Assumptions 2 and 5 hold,
$\beta_{D_{0},D_{1}}=E\left(\frac{V(\Delta D|D_{0},D_{1})}{E\left(V(\Delta D|D_{0},D_{1})\right)}E(S_{2}|D_{0},D_{1})\right),$
and $\beta^{pl}_{D_{0},D_{1}}=0.$
Proof
The proof of the first equality is similar to the proof of the first equality in Theorem 3, and is therefore omitted. Turning to the second equality, it is straightforward to see that
$\beta^{pl}_{D_{0},D_{1}}=$ $\frac{cov(\Delta D_{r,2},\Delta Y_{-1})}{V(\Delta D_{r,2})}$
$=$ $\frac{cov(\Delta D_{r,2},Y_{1}(0)-Y_{0}(0))+cov(\Delta D_{r,2},S_{1}\Delta D_{-1})+cov(\Delta D_{r,2},\Delta S_{-1}D_{0})}{V(\Delta D_{r,2})}$
$=$ $\frac{cov(\Delta D_{r,2},S_{1}D_{1})-cov(\Delta D_{r,2},S_{1}D_{0})}{V(\Delta D_{r,2})}$
$=$ $0$
QED.
5 Application
Reanalysing a two-periods OLS FD regression.
We use the panel data set constructed by Acemoglu et al. (2016), who estimate FD regressions of US industries’ employment evolutions on the evolution of their Chinese imports penetration ratio. The paper’s Table 2 shows several FD 2SLS regressions (Columns (5) to (8)) and one stacked FD OLS regression with two first differences and three periods (Column (2)), but it does not contain any two-periods FD OLS regression. To ensure our empirical application is closely aligned with our theoretical results, we start by revisiting one two-periods FD OLS regression, before considering regressions that were estimated in the paper. We consider the 1999 to 2007 FD regression of US industries employment evolutions on the evolution of their Chinese imports, the two-periods FD regression whose Chinese-imports coefficient is the closest to that in the stacked FD regression in the paper’s Table 2 Column (2). Like in the paper, all linear regressions below are weighted by industries’ 1991 employment, and standard errors are clustered at the three-digit industry level. We find that $\widehat{\beta}=-0.78$ (s.e.=$0.22$), which is very close to the coefficient in the paper’s Table 2 Column (2) ($-0.81$, s.e.=$0.16$). According to this regression, an increase of Chinese imports decreases US employment.
Decomposing $\beta$ under the parallel-trends assumption.
First, we follow de Chaisemartin and D’Haultfœuille (2020), who show that under the parallel-trends assumption, an FD regression estimates a weighted sum of the treatment’s effect in each $(g,t)$ cell, with weights that can be estimated. We estimate those weights, and find that in this specific example, under the parallel-trends assumption $\beta$ estimates a weighted sum of 784 industry×$\times$×year specific effects of Chinese imports on employment, where more than a half of effects are weighted negatively, and where negative weights sum to $-0.35$. Then, the parallel-trends assumption is not sufficient to ensure that $\beta$ is weakly causal. This motivates investigating whether $\beta$ is weakly causal under the assumption that $\Delta D$ is as good as randomly assigned.
Testing if $\Delta D$ and $D_{1}$ are correlated.
We start by testing (7), by regressing $\Delta D$, the change in industries’ import-penetration ratio, on $D_{1}$, industries’ import-penetration ratio in 1999. The coefficient on $D_{1}$ is equal to $0.74$, and it is highly significant (s.e.=$0.16$). Thus, the industries that experienced the largest growth of their imports-penetration ratio from 1999 to 2007 already had a larger imports-penetration ratio in 1999.
Estimating the weights attached to $\beta$ in Theorem 2.
Next, we estimate the weights attached to $\beta$ in Theorem 2. We find that under Assumption 2 and (8), $\beta$ estimates a weighted sum of the average effects of import penetration in 1999 and 2007, where the 2007 average effect receives a weight equal to 1.3, while the 1999 average effect receives a weight equal to -0.3. Thus, under Assumption 2 and (8), the FD regression does not estimate a convex combination of effects, if the effect of import penetration on US employment changes from 1999 to 2007.
Estimation of $\beta_{D_{1}}$.
[…]
Placebo test.
We first conduct a naive placebo test by regressing $\Delta Y_{-1}$, industries employment evolutions from 1991 to 1999, on an intercept and $\Delta D$. The estimated slope of $\Delta D$ is $-0.23$ (s.e. = 0.15), which is statistically insignificant. However, this naive test could be misleading as $\Delta D$ and $\Delta D_{-1}$ are very strongly positively correlated (correlation=0.49). To perform a proper placebo test, using similar arguments as in Section 4.2, we conduct a locally robust version of the test outlined in Section 4.4. Accordingly, we estimate both $E(\Delta D|D_{0},D_{1})$ and $E(\Delta Y_{-1}|D_{0},D_{1})$ using Lasso and run OLS of the corresponding residuals of $\Delta Y_{-1}$ on that of $\Delta D$ [^11] As before, a third degree polynomial in both variables has been specified. The slope coefficient is the estimator of $\beta^{pl}_{D_{0},D_{1}}$. It is equal to $-0.59$ (s.e. = 0.21), [^12] Standard errors were clustered at the three-digit industry level. and thus the data rejects $\beta^{pl}_{D_{0},D_{1}}=0$ at the $5\%$ level. When the previous conditional expectations are estimated by NN, the estimator of the slope coefficient is now $-0.66$ (s.e.=0.21) and then the null $\beta^{pl}_{D_{0},D_{1}}=0$ is similarly rejected. This implies that Assumption 5 is violated: $\Delta D$ does not seem to be as good as random conditional on $(D_{0},D_{1})$.
Re-analysis of the stacked regression in the paper’s Table 2 Column (2).
The paper’s Table 2 Column (2) contains a stacked FD regression with three periods and two first-differences, of $\Delta Y_{t}$ on $\Delta D_{t}$ and a fixed effect (FE) for the second FD. The coefficient on $\Delta D_{t}$ is equal to $-0.81$ (s.e.=$0.16$). Instead, we regress $\Delta Y_{t}-E_{t}(\Delta Y_{t}|D_{t-1})$ on $\Delta D_{t}-E_{t}(\Delta D_{t}|D_{t-1})$, and an FE for the second FD. The conditional expectations in each period are estimated using Lasso. The slope coefficient is equal to $-0.54$, it is insignificant (s.e. = 0.42), and not statistically different from the paper’s coefficient ($H=0.40$). If NN is employed to estimate the conditional expectations, we obtain an estimated coefficient equal to $-0.52$, not different from zero (s.e. = 0.45), and statistically equal to the paper’s coefficient ($H=0.41$).
Re-analysis of the reduced-form of the 2SLS regression in the paper’s Table 2 Column (7).
On top of FD OLS regressions, Acemoglu et al. (2016) also estimate FD 2SLS regressions, where $\Delta D$ is instrumented by $\Delta Z$, the change in industries’ Chinese-import-penetration ratio in eight other high-income countries. We reanalyze the reduced-form regression attached to the two-periods FD 2SLS regression in Table 2 Column (7) of the paper, namely a regression of $\Delta Y$ on $\Delta Z$ taking first differences from 1999 to 2007 as before. The coefficient on $\Delta Z$ is negative and significant ($-1.40$, s.e.=0.38). Instead, we regress $\Delta Y-E(\Delta Y|Z_{1})$ on $\Delta Z-E(\Delta Z|Z_{1})$, where the conditional expectations are estimated using Lasso. The coefficient is smaller than before and insignificant ($-0.86$, s.e.=0.92), and insignificantly different from the paper’s reduced-form coefficient ($H=0.47$). When we reproduce the exercise using NN, the slope coefficient is equal $-0.73$, it is not different from zero (s.e. = 0.91), and it is not significantly different from the paper’s reduced-form coefficient ($H=0.47$).
Table 1: First-difference regressions of employment on Chinese imports
Time periods Dep var Indep var Learner Estimated in paper? Coefficient s.e. N
1999 to 2007 $\Delta Y$ $\Delta D$ None No $-0.78$ $0.22$ 392
1999 to 2007 $\Delta Y-E(\Delta Y|D_{1})$ $\Delta D-E(\Delta D|D_{1})$ Lasso No $-0.60$ $0.39$ 392
1999 to 2007 $\Delta Y-E(\Delta Y|D_{1})$ $\Delta D-E(\Delta D|D_{1})$ NN No $-0.61$ $0.44$ 392
1999 to 2007 $\Delta Y_{-1}$ $\Delta D$ None No $-0.23$ $0.15$ 392
1999 to 2007 $\Delta Y_{-1}-E(\Delta Y_{-1}|D_{0},D_{1})$ $\Delta D-E(\Delta D|D_{0},D_{1})$ Lasso No $-0.59$ $0.21$ 392
1999 to 2007 $\Delta Y_{-1}-E(\Delta Y_{-1}|D_{0},D_{1})$ $\Delta D-E(\Delta D|D_{0},D_{1})$ NN No $-0.66$ $0.21$ 392
Stacked $\Delta Y_{t}$ $\Delta D_{t}$ None Table 2 Column (2) $-0.81$ $0.16$ 784
Stacked $\Delta Y_{t}-E_{t}(\Delta Y_{t}|D_{t-1})$ $\Delta D_{t}-E_{t}(\Delta D_{t}|D_{t-1})$ Lasso No $-0.54$ $0.42$ 784
Stacked $\Delta Y_{t}-E_{t}(\Delta Y_{t}|D_{t-1})$ $\Delta D_{t}-E_{t}(\Delta D_{t}|D_{t-1})$ NN No $-0.52$ $0.45$ 784
1999 to 2007 $\Delta Y$ $\Delta Z$ None RF of Table 2 Column (7) $-1.40$ $0.38$ 392
1999 to 2007 $\Delta Y-E(\Delta Y|Z_{1})$ $\Delta Z-E(\Delta Z|Z_{1})$ Lasso No $-0.86$ $0.92$ 392
1999 to 2007 $\Delta Y-E(\Delta Y|Z_{1})$ $\Delta Z-E(\Delta Z|Z_{1})$ NN No $-0.73$ $0.91$ 392
The table shows regressions of US industries employment evolutions $\Delta Y$ on the evolution of their Chinese import-penetration ratio $\Delta D$, using the data of Acemoglu et al. (2016). Regressions are weighted by industries’ 1991 employment, and standard errors are clustered on 135 three-digit industries, as in the paper. The first line shows a FD regression from 1999 to 2007. The second and third lines are similar, but control for $E(\Delta D|D_{1})$ using a locally-robust estimator, where conditional expectations are estimated by Lasso and NN, respectively. The fourth line performs a naive placebo test. The fifth and sixth lines conduct placebo tests controlling for $E(\Delta D|D_{0},D_{1})$, where conditional expectations are estimated by Lasso and NN, respectively. The seventh line re-estimates the stacked regression in the paper’s Table 2 Column (2). The eighth and ninth lines are similar to the seventh, but control for $E_{t}(\Delta D_{t}|D_{t-1})$ using a locally-robust estimator, where conditional expectations are estimated by Lasso and NN, respectively. The tenth line re-estimates the reduced-form of the paper’s 2SLS regression in Table 2 Column (7). The last two lines are similar, but control for $E(\Delta Y|Z_{1})$ using a locally-robust estimator, where conditional expectations are estimated by Lasso and NN, respectively.
6 Conclusion
This paper’s main message is that under a causal model in levels, first-difference regressions may suffer from an omitted variable bias if the treatment change $\Delta D$ is correlated to the baseline treatment $D_{1}$, and the treatment effect changes over time. A simple fix is to control for $E(\Delta D|D_{1})$.